### v1

# Accounting for motion of supernova host galaxy in statistical inference from SNIa data

Abstract
We investigate the impact of peculiar motion of Type Ia supernova host galaxies on cosmological parameter estimation. This motion causes their redshift to deviate from that of the comoving observer at their position and is a source of noise. To this end, we develop an estimator for parameter estimation in models with errors in independent variables. Using the Bayesian framework, errors in independent variables are treated as nuisance parameters by making the independent variables parameters of the model. Our method applied to the Pantheon sample of Type Ia supernova indicates a few percent shift in the central values of inferred cosmological parameters. For the $w$CDM model, we find that accounting for peculiar velocities makes the data marginally more consistent with the cosmological constant model. By using simulated data, we show that not accounting for peculiar velocities will significantly impact parameter estimation from higher precision future data sets.
1 Introduction
The spatially flat $\Lambda$CDM model, based on the cosmological principle of homogeneity and isotropy, constituting cold dark matter and the cosmological constant that dominates the universe at present, has become the standard description of our observed universe. It has been remarkably successful in describing the observed anisotropies in the Cosmic Microwave Background (CMB)[1], large-scale clustering of matter [2, 3, 4], and the background expansion history inferred from the redshift and brightness measurements of Type Ia supernovae (SNIa) [5, 6, 7]. In addition, a growing number of space and earth missions have ushered in the era of precision cosmology [8], which allows us to constrain cosmological parameters at a sub-percent level of precision. For example, the measurement of the CMB temperature and polarization anisotropies by the Planck mission has allowed us to constrain the Hubble constant within an uncertainty of $0.5$ km/s/Mpc [1]. This increasing quality and quantity of data has enabled us to extract more detailed information about the present and past of the universe.
However, the onset of the era of greater precision has also revealed systematic discrepancies in the standard model of cosmology, such as the observed tensions in $H_{0}$ and $S_{8}$ [9, 10, 11]. Taking these discrepancies as a hint of new physics beyond the standard $\Lambda$CDM model, several modifications to the theory have been proposed in the literature, on the contrary, untreated or systematic errors in the analysis as a source of observed discrepancies remain a less explored but plausible alternative. For example, there are certain anomalies in the CMB observations, such as lensing anomalies, whose presence depends on the method used for statistical analysis. For example, the standard Plik and Commander likelihood of Planck 2018 data prefer a higher value of the lensing parameter $A_{L}$ than the prediction of the $\Lambda\rm CDM$ model, $A_{L}=1$. However, in the likelihood NPIPE based on the same Planck data but different modeling of systematics, there is no such anomaly [12]. It was shown in [13] that alleviating these tensions with extensions to the $\Lambda$CDM model is not sufficient to claim a hint for physics beyond the $\Lambda$CDM model. This also highlights the fact that increasing the precision in measurement also requires more accurate statistical analysis techniques to obtain consistent results from different observational probes.
Although highly precise instrumentation equipped with artificial intelligence and machine learning is now widely used in astronomy, astronomical observations are never free from errors. These errors can originate from instrumental limitations, atmospheric interference (for ground-based telescopes), or cosmic background noise. Furthermore, systematic biases introduced during data processing and limitations of observational pipelines contribute to these inaccuracies. Despite this, while fitting a model to the data, it is very common to consider errors only on the main observable (dependent variable), treating subsidiary observables (independent variables) as perfectly measured. In part, this is due to the fact that the usual statistical techniques are based on the likelihood function that is derived from the errors in the main observable. This is usually justified if the errors in the subsidiary observables are small compared to the errors in the dependent variable.
In statistics literature, the study of model fitting when there are errors in both dependent and independent variables is called errors-in-variables models [14]. Many different methods have been proposed to deal with this issue if additional information can be obtained about the independent variable. The instrumental variable method [14] considers another variable that is related to the true regressor with some known relation. The method of repeated measurement [14] uses multiple observations of the same variable to minimize error. For the special case, when the model is a monotonic and differentiable function of the regressor variable, the likelihood can be written by inverting the model function [15]. However, there is no general method for fitting arbitrary errors-in-variable models.
In the context of cosmology, this issue could arise on various occasions. For example, it becomes relevant in estimating the cosmological parameters from the magnitude–redshift data of SNIa. Usually, the redshift determination of these supernovae is considered to be free of errors. There are at least two sources of errors in the determination of the redshift. Spectroscopic redshifts can be determined with a relative accuracy better than $10^{-4}$. The second source of error arises from the line-of-sight peculiar velocities owing to gravitational interaction [^1] Cosmological peculiar velocities are a key prediction of the gravitational collapse model of structure formation. Their impact can be detected in the redshift-space two-point correlation function of galaxies (e.g. [16] and the references therein for details). These velocities can also be detected statistically using the kSZ (Kinetic Sunayaev-Zeldovich) effect (e.g. [17, 18]). The line-of-sight peculiar motion of host galaxies of supernovae contributes to the total observed redshift and cannot be disentangled from the contribution from cosmological expansion [^2] The contribution of peculiar motion also arises from our motion with respect to the cosmic rest frame. The magnitude and direction of this component can be determined from the measurement of the CMB dipole. It is corrected for in the available SNIa data as in the Pantheon sample.. From linear perturbation theory, the RMS of peculiar velocity varies from a few hundred km/sec to a few thousand km/sec for scales 1 Mpc to 100 Mpc. The modeling of the peculiar motion of individual objects requires taking into account nonlinear effects, which could further enhance this effect. The impact of peculiar velocities could be more than an order of magnitude larger than the instrumental uncertainty but still small compared to Hubble recessional velocities except in the nearby universe. The impact of peculiar motion on the cosmological parameters has been studied in the literature using N-Body simulations [19, 20, 21, 22]. The peculiar velocity correction in the Pantheon+ SNIa sample was obtained in [23]. These studies focus mainly on the determination of the Hubble constant.
In this work, we analyze the impact of peculiar velocities on the determination of cosmological parameters using SNIa data by studying the problem of fitting models with errors in both dependent and independent variables. We analyze the Pantheon SNIa sample [24] using three estimators—one without peculiar motions, one based on a locally linear magnitude–redshift relationship, and finally, one that accounts for the theoretical model without any approximation. We then apply our estimators on simulated data with parameters (e.g. number of SNIa, errors on magnitude) typical of the current and future SNIa data sets. For our study, we consider two cosmological models: a spatially flat $\Lambda$CDM model and a $w$CDM model with constant $w$.
This paper is organized as follows. In Section 2, we discuss in general our method for dealing with the problem of fitting arbitrary models with errors in both variables. In Section 3, we present the magnitude–redshift relation as an errors-in-variables model. In Section 4, we apply this method to estimate cosmological parameters from supernova data for the $\Lambda$CDM and the $w$CDM model, and discuss the modification in the results as compared to the standard model fitting method that does not consider the contribution of peculiar motion. In Section 5, we conclude with a discussion of the results and potential future applications of our method. In Appendix A, we derive an expression for likelihood in the linear approximation of the magnitude–redshift relation.
2 Methodology
The problem of regression with an arbitrary errors-in-variables model is interesting and, to the best of our knowledge, not a commonly addressed problem in astronomy. Since astronomical data usually have errors in independent variables, this could have general applications in a variety of problems in astronomy. In this section, we describe a Bayesian approach to this problem. In Section 2.1, we describe our method in general while motivating our choice of the particular form of estimator that we propose. Then, in Section 3, we specialize it for the estimation of cosmological parameters from the magnitude–redshift data of the Pantheon sample of supernovae [24]. We consider the contribution of peculiar motion of supernova host galaxies to the redshift uncertainty, which leads to its formulation as an errors-in-variables problem. We also compare it to an approach that uses a local linear approximation of the magnitude–redshift relation and uses known methods of handling errors-in-variables for linear models.
2.1 Estimator for Errors-in-Variables
Quite often, the output of experiments is a pair $(x_{i},y_{i})$ [^3] It is worth noting that in general, the data could have several independent variables $(x_{i},y_{i},z_{i},\cdots)$. It is straightforward to generalize this analysis to such multivariate cases. of $N$ data points with uncertainties, assuming the randomness to be Gaussian, $\sigma_{x_{i}}$ and $\sigma_{y_{i}}$, associated with each observed value $x_{i}$ and $y_{i}$ respectively. The quantity $y_{i}$ often depends on the independently measured quantity $x_{i}$ through a model that could be given through a mathematical relation
$y_{i}=f(\theta,x_{i}),$ (2.1)
or, in general, could be given by an algorithm such as a numerical simulation or a complex mathematical operation. Here, $f(\theta,x_{i})$ is a function of the $m$ model parameters, $\theta\equiv(\theta_{1},\theta_{2},...\theta_{m})$ and the variable $x_{i}$. In our notation, if any expression depends explicitly on a single variable, then we notate it with a subscript; else, we use unscripted symbols to refer to them collectively. We shall soon focus on the case of apparent magnitude versus redshift data that comprises the measured pair $(z_{i},m_{i})$, which might be helpful to the reader while reading the following discussion to make it less abstract.
If the function $f(\theta,x_{j})$ is a linear function of $x_{j}$, then estimating the parameters $\theta$ from data that have errors on both axes is quite straightforward (see, for example, [25] and the derivation in Appendix A). This method is based on the conversion of the $x$ error into the $y$ error through the slope of the linear function and adding it in quadrature to the $y$ error while constructing the likelihood function. However, when the model is nonlinear in $x_{j}$, this trick works only by approximating the fitting function as locally a linear function, as shown for the special case of the magnitude–redshift relation in the Appendix A. Although, in general, a nonlinear model could refer to nonlinear dependence on $\theta$ or $x_{i}$, in this paper, by nonlinear model, we always refer to the latter case.
2.1.1 Estimator for nonlinear models
To keep things general, let $P_{\rm Y}(\epsilon_{Y})$ be the probability distribution function for random statistical noise in the measurement of $y$, so that the observed values are offset from their true value by a random number, $y_{i}=f(\theta,x_{i})+\epsilon_{Yi}$, where $\epsilon_{Yi}$ is the random error in the measurement $y_{i}$. Therefore, the likelihood function is given by
$\mathcal{L}(\mathcal{D}\mid\theta)=\prod_{i}P_{Y}(y_{i}-f(\theta,x_{i})),$ (2.2)
where $\mathcal{D}$ collectively refers to the pair $(x,y)$. In Bayesian statistics, using Bayes’ theorem, this likelihood function can be used to write the posterior probability distribution for parameters $\theta$ as
$P(\theta\mid\mathcal{D};I)\propto\mathcal{L}(\mathcal{D}\mid\theta;I)P(\theta\mid I)$ (2.3)
where $P(\theta\mid I)$ is the prior probability distribution function for $\theta$, and the addition of $I$ in various terms symbolizes all the prior information that we may have (for details, see, e.g., [26]). Until now, the variables $x$ have been considered given and fixed. To incorporate error in variables $x$ in the posterior distribution, we now turn them into undetermined parameters $x^{*}$, equivalent to parameters $\theta$, in terms of which the likelihood function can be written as
$\mathcal{L}(\mathcal{D}\mid\theta,x^{*};I)=\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))$ (2.4)
Note that conversion of $x$ to parameter means that now the data $\mathcal{D}$ refers only to $y$. The information about errors in $x$ and their measured values can be brought into our analysis as a prior on $x^{*}$ through [^4] We can derive the same expressions by noting that the joint probability of errors $\epsilon_{X}$ and $\epsilon_{Y}$ is given by $P_{X}(\epsilon_{X})P_{Y}(\epsilon_{Y})$, thus the errors on $x$ can be considered to be a part of likelihood of a data set comprising the pair $(x,y)$
$P(x^{*}_{i}\mid I)=P_{\rm X}(x^{*}_{i}-x_{i}),$
where $P_{\rm X}$ is the probability distribution for errors in variable $x$ (assumed for simplicity to be the same for all data points, but this can be easily generalized). The modified posterior distribution is given by
$P(\theta,x^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))P(x^{*}_{i}\mid I)P(\theta\mid I)$ (2.5)
If the random noise is Gaussian, then the log of the likelihood function is given by
$\ln\mathcal{L}(\mathcal{D}\mid\theta,x^{*})=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{y_{i}-f(\theta,x^{*}_{i})}{\sigma_{y_{i}}}\right]^{2}\equiv-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2},$ (2.6)
and the log of the prior probability distribution function for $x$ is given by
$\ln{P(x^{*}_{i}\mid I)}=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{x^{*}_{i}-x_{i}}{\sigma_{x_{i}}}\right]^{2}\equiv-\frac{\chi^{2}_{X}(x^{*})}{2}$ (2.7)
In terms of these variables, the posterior distribution can be written as
$P(\theta,x^{*}\mid\mathcal{D})\propto\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)$ (2.8)
We can marginalize $x^{*}$ to obtain the probability distribution for $\theta$
$P(\theta\mid\mathcal{D})\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)$ (2.9)
In practice, marginalization is more conveniently carried out through MCMC, as we discuss later.
Since the total number of parameters now exceeds the number of data points, there is, in principle, the possibility of overfitting the data. In order to obviate this possibility, we add a regularization term as a prior probability that our physical model should be such that the mean value of $\chi_{Y}^{2}$ is close to degrees of freedom of the model [^5] This expectation is technically true only for linear models, but holds quite well in general for small errors as models can then be roughly treated as locally linear.. For the case of Gaussian noise in the dependent variable, this is just the probability distribution function for $\chi_{Y}^{2}$ given by
$P_{\chi^{2}}(\chi^{2}_{Y})=\frac{\chi_{Y}^{k-2}\exp(-\chi^{2}_{Y}/2)}{2^{k/2}\Gamma(k/2)},$ (2.10)
where $k=N-m$ is the desired degrees of freedom of the fit (not accounting for noise in the $x$ values). For example, for the case of fitting luminosity distance to distant supernovae, this would be the total number of supernovae minus the number of cosmological parameters being fitted. Therefore, our final form for the posterior distribution function for $\theta$ can be obtained by multiplying it with the probability distribution function for $\chi_{Y}^{2}$
$P(\theta\mid\mathcal{D})\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)P_{\chi^{2}}(\chi_{Y}^{2})\,.$ (2.11)
Regularization is commonly used in fitting complex models with a large number of parameters as it acts as a constraint and facilitates stable maximization of the posterior distribution (e.g. [26]). However, as we note later, the addition or omission of this term may not always change the dimensionality of the fit, as explained further in the context of the specific cosmological application of this method presented in the paper.
3 Magnitude–Redshift relation
Although the measurement error in the observed redshift of a supernova host galaxy is negligibly small, due to the possible peculiar motion of the host galaxy with respect to an observer at rest with respect to the universe, their observed redshift is shifted with respect to the theoretical cosmological redshift that encodes the expansion history of the universe. To estimate this shift, we note that
$\frac{\nu_{0}}{\nu_{\rm obs}}=\frac{\nu_{0}}{\nu_{\rm com}}\frac{\nu_{\rm com}}{\nu_{\rm obs}}\approx(1+z)(1+{v_{p}}/{c})\approx 1+z+{v_{p}}/{c}$ (3.1)
where $\nu_{\rm obs},\nu_{0}$, and $\nu_{\rm com}$ are the frequencies observed at earth, in the rest frame of the galaxy and a comoving observer at the position of the galaxy, and $v_{p}$ is the component of velocity of a galaxy with respect to the comoving observer parallel to the line-of-sight from the observer to the supernova host galaxy. The approximation in the last equation assumes $v_{p}/c\ll 1$, so we can neglect the transverse Doppler effect. For small redshifts where this correction dominates, it suffices to ignore the $zv_{p}/c$ term, leading to the final form.
Peculiar motions also cause beaming of light towards the observer, alter the observed flux, and cause additional k-correction (e.g. see [27] for the combined impact of all these effects on the number count). The first two effects act in concert, as an object moving towards an observer would appear closer and less redshifted. The standard analysis already takes into account measurement errors in the luminosity of an SNIa, and these errors are much larger than what we expect from the peculiar motion of host galaxies. However, this is not true for redshift measurement errors that, as already noted, are small compared to peculiar motions. At low redshifts, this effect is enhanced as peculiar redshift/blueshift can become comparable to the redshift of the host galaxy. As more and more supernovae are observed, the effective statistical error in luminosity distance reduces at any redshift due to statistical averaging. Thus, we need techniques that allow us to handle errors on redshift, even if they are small.
The luminosity distance versus redshift relation is usually expressed in terms of the apparent magnitude of distant supernovae. To apply the analysis given in the previous section, we set the redshift $z_{i}$ equal to $x_{i}$ and the apparent magnitude $m_{i}$ to the function $f(\theta,x_{i})$. The apparent magnitude $m$ and the luminosity distance $D_{\rm L}$ of a supernova of known absolute magnitude $M$ are related through
$m=M+5\log\left(\frac{D_{\rm L}}{\text{Mpc}}\right)+25\;.$ (3.2)
For a flat Friedmann–Lemaître–Robertson–Walker universe, the luminosity distance of a source at redshift $z$ can be expressed as
$D_{\rm L}=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{h(z^{\prime})},$ (3.3)
where the dimensionless Hubble parameter is given by
$h(z)=\left[\Omega_{\rm M}(1+z)^{3}+\Omega_{\rm DE}(1+z)^{3(1+w)}\right]^{1/2}\,,$ (3.4)
where $\Omega_{\rm M}$ and $\Omega_{\rm DE}$ are the present-day matter and dark energy densities respectively. The dark energy and pressure are related through the parameter $w$ through $p_{\rm DE}=w\rho_{\rm DE}$. If dark energy is assumed to be the cosmological constant, then $w=-1$ in this parametrization. The apparent magnitude can be written as
$m=M+5\log(cH^{-1}_{0}/{\rm Mpc})+5\log\left[(1+z)\int_{0}^{z}\frac{dz^{\prime}}{h(z)}\right]+25\,.$ (3.5)
Here, $w$, $\Omega_{\rm M}$, and $H_{0}$ are free parameters of the model. However, note that $H_{0}$ is degenerate with $M$, which allows us to estimate only their combination.
4 Results
In cosmological studies, our primary concern is with the parameters $w$ and $\Omega_{\rm M}$. For presenting our results, we need to marginalize various parameters, which means integrating them in the posterior distribution. Also note that in Eq. (2.11), we need to integrate over $x^{*}$, which corresponds to the redshift $z$ in this case. Clearly, it is not feasible to integrate over so many variables. Therefore, to carry out marginalization, we employ the Markov Chain Monte Carlo (MCMC) sampling technique. Specifically, we develop and implement the Metropolis-Hastings algorithm within the MCMC framework to draw samples from the posterior distribution and use GetDist [28] python library for the analysis of MCMC samples.
Figure 1: The figure displays one-dimensional posterior distributions and two-dimensional (1- and 2-$\sigma$) contours of $\Lambda$CDM parameters from Pantheon data using three different estimators. Estimator $\mathcal{E}_{1}$ (red) ignores any contribution to redshift from the peculiar motion of supernovae host galaxies, $\mathcal{E}_{2}$ considers it for the linearized model (see Appendix A for details), while $\mathcal{E}_{3}$ (green) represents our estimator, which incorporates the contribution from peculiar motion for the exact model.
Parameters Priors Best-fit values
[…]
$\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.286\pm 0.012$ $0.296\pm 0.014$ $0.302\pm 0.016$
$H_{0}$ $\mathcal{U}(60,80)$ $73.59\pm 0.229$ $73.34\pm 0.287$ $73.13\pm 0.343$
[…]
Although distant SNIa are not ideal for determining $H_{0}$ due to degeneracy with $M$, in the following sections, we also present results for the Hubble parameter for the $\Lambda$CDM model. In order to break the aforementioned degeneracy, we use the absolute magnitude $M=-19.2478$ corresponding to the SH0ES [29] value of $H_{0}=73.04$ km/s/Mpc for the Pantheon sample.
For the redshift error, we assume that each observed $z_{i}$ in the data deviates from its true value $z^{*}_{i}$, which falls within a specified uncertainty range $z^{*}_{i}\in[z_{i}^{\rm min},z_{i}^{\rm max}]$, where the min and max values are chosen to correspond to our assumptions about peculiar velocities. This is imposed even if we assume the distribution of redshift to be nonuniform. From analytical theory or numerical simulations, we can estimate the variance of galaxy peculiar velocities as a function of redshift. Also, note that only the line-of-sight component of the peculiar velocity contributes to the change in redshift. By the isotropy of the peculiar velocity dispersion, we have $\sigma_{v_{\rm 3D}}=\sqrt{3}\sigma_{v_{p}}$
In our analysis below, we consider three distinct cases to present our results. The standard case is labeled as $\mathcal{E}_{1}$, where we do not account for the redshift error. In Appendix A, we derive an estimator based on the linearized magnitude–redshift relation labeled $\mathcal{E}_{2}$ to include the redshift error using the likelihood function (Eq. (A.6)). This is of interest since the MCMC-based final estimator ($\mathcal{E}_{3}$) using the method described in Section 2.1.1 (Eqs. (2.2)–(2.11)), which includes the contribution of peculiar velocity to the observed redshift by treating each redshift as a parameter, can be computationally expensive for large data sets.
4.1 Application to Pantheon Data
We apply our method to fit the $\Lambda$CDM and the $w$CDM models to the magnitude–redshift data of Type Ia supernovae in the Pantheon sample [24]. This data set contains 1048 spectroscopically confirmed SNIa. For all cases, we run 12 independent MCMC chains for $10^{6}$ steps. We use the Gelman-Rubin criterion of convergence. For the $\Lambda$CDM model, we achieve $R-1<0.001$, while for the $w$CDM model it is $R-1<0.005$. For $z$-errors, we use uniform priors for $v_{pi}$ in the range: $-1000\leq v_{pi}\leq 1000\,\rm km\,s^{-1}$ around the measured redshift. To verify convergence and test its rate, we increase the range of peculiar velocities to $-3000\leq v_{pi}\leq 3000\,\rm km\,s^{-1}$. Increasing this interval reduces the acceptance rate and increases $R-1$ for the fixed number of MCMC steps. However, $R-1$ is always less than $0.01$. Thus, it leads to slower convergence, but our results remain unchanged; we also obtain the same results by assuming a Gaussian distribution for peculiar velocities, imposing only the condition that $x^{*}>0$.
Figure 2: The figure shows posterior distributions and contours of $w$CDM parameters from Pantheon data using three different estimators. It follows the convention used in Figure 1.
Parameters Priors Best-fit values
[…]
$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.6)$ $0.348\pm 0.035$ $0.335\pm 0.043$ $0.330\pm 0.046$
$H_{0}$ $\mathcal{U}(60,80)$ $74.07\pm 0.368$ $73.80\pm 0.546$ $73.58\pm 0.703$
$w$ $\mathcal{U}(-2,0)$ $-1.24\pm 0.14$ $-1.17\pm 0.17$ $-1.14\pm 0.18$
[…]
It should be noted that our results are insensitive to regularization (Eq. (2.11)). The regularization term only changes the acceptance rate, indicating that the inclusion of redshift errors does not increase the dimensionality of our fitted physical model. This is not surprising if we note that overfitting requires the fitting function to be able to pass through most of the noisy data points. This would require the fitting function to have the flexibility that Eq. (3.2) does not have, even if we convert the redshift to a variable. The function remains a monotonic function of redshift and thus does not cause overfitting of the data.
Figures 1 and 2 display the posterior distribution and the confidence regions $1\sigma$ and $2\sigma$ for the $\Lambda$CDM and $w$CDM models for the Pantheon data. Different colors represent the three different estimators used. The red color represents the standard case, $\mathcal{E}_{1}$, blue represents the linearized estimator, $\mathcal{E}_{2}$, and the green represents the MCMC estimator, $\mathcal{E}_{3}$. In both cosmological models, we have assumed uniform priors for the cosmological parameters, which are listed in Tables 1 and 2. Perusal of Figures 1 & 2 indicates that accounting for peculiar velocities in the current data does not significantly change the estimated parameters. However, note that in Table 2 for the $w$CDM model, the marginalized estimate for the dark energy parameter $w$ is found to be marginally more consistent with the $\Lambda$CDM model, indicating that even at the current level of precision, the data could indicate non-trivial dark energy if systematic errors in redshift are not corrected. As the number of supernovae increases, this correction is likely to become even more significant. Next, we consider simulated data to investigate this issue further.
Figure 3: Constraints on the $\Lambda$CDM parameters from synthetic data with magnitude error $\Delta m=0.2$ (left) and $\Delta m=0.05$ (right) The colour scheme of Figure 1 is followed.
Models Parameters Priors True values Best-fit values
[…]
$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.217\pm 0.014$ $0.240\pm 0.016$ $0.280\pm 0.020$
$H_{0}$ $\mathcal{U}(60,80)$ $73.0$ $75.60\pm 0.338$ $74.88\pm 0.391$ $73.45\pm 0.482$
$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.368\pm 0.023$ $0.356\pm 0.028$ $0.320\pm 0.048$
$H_{0}$ $\mathcal{U}(60,80)$ $73.0$ $78.01\pm 0.642$ $77.09\pm 0.849$ $74.29\pm 1.024$
$w$ $\mathcal{U}(-3,0)$ $-1.0$ $-1.89\pm 0.213$ $-1.68\pm 0.236$ $-1.22\pm 0.228$
Table 3: Priors and best-fit values for the $\Lambda$CDM and the $w$CDM parameters obtained from synthetic data with $\Delta m=0.2$.
4.2 Application to Synthetic Data
For simulations, we first consider the parameters of the Pantheon sample: 1048 supernovae in the redshift range $z\in[0.01,2.26]$ with magnitude errors, $\Delta m\simeq 0.2$ (for simulation, we assume Gaussian random errors, $\mathcal{N}(0,0.20)$). We also consider a future SNIa project with errors four times smaller in magnitude, $\mathcal{N}(0,0.05)$. This is possible with a sample 16 times larger than the Pantheon sample with the magnitude errors obtained from binning the data in redshift (this might be achieved with Zwicky Transient Facility (ZTF)[30] and the Large Synoptic Sky Survey Telescope (LSST) [31, 32], albeit with higher supernovae density at $z<0.5$, see, e.g. [33] for details). The cosmological and other parameters are fixed to the following values: $\Omega_{\rm M}=0.30$, $H_{0}=73~{}\rm km~{}s^{-1}~{}Mpc^{-1}$, $w=-1.0$, and the absolute magnitude of supernova, $M=-19.24$. To account for peculiar velocities, each supernova is assigned an error on its redshift, $\mathcal{U}(0,0.003)$, drawn from a uniform distribution.
Our main results are shown in Figures 3 and 4 together with Tables 3 and 4. A comparison of the analysis of synthetic data with Figures 1 and 2 enables us to discern the following: (a) there is bias in the estimation of cosmological parameters using estimators ${\mathcal{E}}_{1}$ and ${\mathcal{E}}_{2}$. The bias is larger if the peculiar velocity contribution is totally neglected (${\mathcal{E}}_{1}$) and is partly alleviated if the linear approximation (Appendix A) is used and (b) while the overlapping confidence regions for $\Delta m=0.2$ do not allow us to determine this bias unambiguously using the current data, the analysis of the future data without taking into account peculiar velocities of host galaxies will result in significant bias in the determination of cosmological parameters. In particular, the analysis of SNIa data without accounting for peculiar velocities could bias the nature of dark energy, an important issue from the point of view of both cosmology and theoretical physics, and much debated in recent times following DESI results (e.g. [4]).
Figure 4: Constraints on the $w$CDM parameters from synthetic data with $\Delta m=0.2$ (left) and $\Delta m=0.05$ (right).
Models Parameters Priors True values Best-fit values
[…]
$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.30$ $0.233\pm 0.005$ $0.280\pm 0.005$ $0.298\pm 0.005$
$H_{0}$ $\mathcal{U}(60,80)$ $73.04$ $75.42\pm 0.19$ $74.03\pm 0.129$ $73.22\pm 0.118$
$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.30$ $0.38\pm 0.010$ $0.322\pm 0.012$ $0.310\pm 0.015$
$H_{0}$ $\mathcal{U}(60,80)$ $73.04$ $77.74\pm 0.252$ $74.86\pm 0.304$ $73.41\pm 0.243$
$w$ $\mathcal{U}(-3,0)$ $-1.0$ $-1.87\pm 0.08$ $-1.20\pm 0.07$ $-1.05\pm 0.06$
Table 4: Priors and best-fit values for the $\Lambda$CDM and the $w$CDM parameters obtained from synthetic data with $\Delta m=0.05$.
5 Conclusion and Discussion
Cosmology has become a precision science in the past few decades. This has been achieved by an unprecedented increase in both the quality and quantity of data—Planck CMB temperature and polarization maps, Large galaxy surveys such as SDSS, and the detection of thousands of SNIa, etc. This also means physical effects hitherto considered beyond the ability of astronomical instruments are well within reach. However, such rapid progress also requires one to constantly improve existing statistical techniques and develop new ones. This paper is an effort in that direction.
The aim of the paper is two-fold. First, we develop a general Bayesian method to analyse data with errors in both independent and dependent variables (discussed in detail in section 2.1.1). Second, we employ this statistical estimator to SNIa data to test its applicability, efficacy, and relevance to the current and future SNIa and other cosmological data sets. In particular, our aim is to discern the role of peculiar velocities of host galaxies of SNIa on the determination of cosmological parameters.
In Figures 1 and 2, we apply our proposed method on the Pantheon SNIa data for two cosmological models we employ for illustration. The figures show that the current data do not require us to account for the nonzero peculiar velocity of the SNIa host galaxies. We next apply our estimators on synthetic data, which partly confirms our results with Pantheon data, for parameters of the current data (Figures 3 and 4). However, we show that the future data (e.g. Zwicky Transient Facility (ZTF)[30] and the Large Synoptic Sky Survey Telescope (LSST) [31, 32]) will be very sensitive to peculiar velocities of SNIa host galaxies, and we would need appropriate statistical tools (section 2.1.1) to extract cosmological parameters from such a data set without bias (Figures 3 and 4).
One possible improvement of our work is more detailed modelling of peculiar velocities of host galaxies. It is difficult to predict the peculiar velocities of host galaxies without N-body simulations (e.g. [19] for a recent effort on the impact of peculiar velocities on Hubble’s constant measurement). This velocity would also depend on the location of host galaxies, e.g., a host galaxy in a rich cluster of galaxies as opposed to an isolated galaxy. In addition, and importantly, velocities are correlated across 100s of Mpc (e.g. [34] for details and further references), and this correlated component needs to be accounted for in a more detailed discussion (e.g. [33]). Also, the aim of our paper is to treat peculiar velocities as nuisance parameters and to determine their impact on the interpretation of SNIa data. However, it is clear from the context that one might seek to infer cosmological peculiar velocities from the SNIa data. From cosmological data, peculiar flows are difficult to measure for individual objects and their presence is revealed through redshift-space distortion of galaxy correlation functions (e.g. [34] for details) and kSZ effect of clusters (e.g. [17, 18] for further details). Our proposed method might complement this ongoing effort to measure cosmological peculiar flows.
In future work, our aim is to extend our work with a better modeling of peculiar velocities along with studying the feasibility of their detection using the SNIa data. In addition, a combined analysis using SNIa data sets Pantheon$+$, DESY5, and Union3 might improve our results, which we hope to investigate in a later work. We also expect that our work will find general applicability in the analysis of other cosmological data sets, which we might study in upcoming works.
Acknowledgments
UU would like to acknowledge Yashi Tiwari for her valuable input throughout the project, particularly her insightful feedback on improving the efficiency of our Python code. UU also thanks Somnath Bharadwaj, Purba Mukherjee, and Avinash Paladi for their helpful comments.
Appendix A Linear Approximation
The problem of fitting errors-in-variable models is quite straightforward for linear models. This is because the linear model allows us to marginalize the likelihood over the true regressor analytically, resulting in a very simple likelihood that depends only on the observed values of the variables [25]. Since our model is not linear, we cannot apply it exactly. We compare our method with the standard method applied to the linear approximation of the model. Here, we derive an expression for the likelihood of the linear approximation of our model.
Apparent magnitude as a function of redshift, $z$, is given by
$m(z)=M+5\;\log\left(\frac{D_{\rm L}(z)}{\text{Mpc}}\right)+25\;,$ (A.1)
where $D_{\rm L}(z)$ is the luminosity distance given by,
$D_{\rm L}(z)=(1+z)\int_{0}^{z}\frac{c\;dz}{H(z)}.$ (A.2)
Considering errors in the measurement of both the apparent magnitude, $m$, and redshift, $z$, the likelihood can be expressed as
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z^{*}_{i})}{\sigma_{m_{i}}}\right)^{2}-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{z_{i}-z^{*}_{i}}{\sigma_{z_{i}}}\right)^{2}$ (A.3)
This likelihood cannot be used to constrain the parameters of the model from the data since we do not have the true redshift values $z_{*}$, only the observed values $z$. Taylor expanding the apparent magnitude about the observed redshift and keeping terms up to linear order in redshift, we can write
$m(z^{*})=m(z)+m^{\prime}(z)(z^{*}-z)$ (A.4)
where $m^{\prime}(z)$ is the derivative of $m(z^{*})$ w.r.t. $z^{*}$ evaluated at $z^{*}=z$, and is given by
$m^{\prime}(z)=\frac{5}{\ln 10}\left[\frac{1}{1+z}+\frac{1}{H(z)\int_{0}^{z}\frac{dz^{*}}{H(z^{*})}}\right]$ (A.5)
With the above approximation, eq. (A.4), the likelihood takes the form,
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z_{i})}{\sigma_{m_{i}}}\right)^{2}-\\ \frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{m^{\prime 2}(z_{i})}{\sigma^{2}_{m_{i}}}+\frac{1}{\sigma^{2}_{z_{i}}}\right)(z^{*}_{i}-z_{i})^{2}-\frac{m^{\prime}(m_{i}-m(z_{i}))}{\sigma^{2}_{m_{i}}}(z^{*}_{i}-z_{i})\right]$
Integrating over the true but unknown redshift $z^{*}$ using the standard Gaussian integral, we obtain the marginalized likelihood,
$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\frac{\left(m_{i}-m(z_{i})\right)^{2}}{\sigma^{2}_{m_{i}}+m^{\prime 2}(z_{i})\;\sigma^{2}_{z_{i}}}$ (A.6)
Now, the likelihood depends only on the observed data. If we do not know the values of $\sigma_{z_{i}}$, as in our case, we can fix it at some value or make it a parameter. In this work, we chose to fix it at $\sigma_{z}=0.003\;$, which corresponds to the peculiar velocity of $1000$ km/s.

### v2

@@
 Abstract
-We investigate the impact of peculiar motion of Type Ia supernova host galaxies on cosmological parameter estimation.
-This motion causes their redshift to deviate from that of the comoving observer at their position and is a source of noise.
-To this end, we develop an estimator for parameter estimation in models with errors in independent variables.
-Using the Bayesian framework, errors in independent variables are treated as nuisance parameters by making the independent variables parameters of the model.
-Our method applied to the Pantheon sample of Type Ia supernova indicates a few percent shift in the central values of inferred cosmological parameters.
-For the $w$CDM model, we find that accounting for peculiar velocities makes the data marginally more consistent with the cosmological constant model.
-By using simulated data, we show that not accounting for peculiar velocities will significantly impact parameter estimation from higher precision future data sets.
-1 Introduction
+We introduce a Bayesian method to estimate peculiar velocities of Type Ia supernova (SNIa) host galaxies by employing the magnitude-redshift relationship of SNIa.
+Random peculiar motions act as noise in the estimation of redshift, and constitute independent variables in the SNIa data.
+We develop a method to take into account errors in independent variables for general nonlinear models.
+Using the MCMC sampling technique, we implement numerical codes to estimate peculiar velocities along with other cosmological parameters.
+In this paper, we study the impact of these velocities on cosmological parameters by treating them as nuisance parameters.
+We apply our proposed method on the Pantheon sample of SNIa and show a few percent shift on the central values of inferred cosmological parameters.
+Although the current data are not particularly sensitive to this error, using simulated data, we also gauge the efficacy of our method on the future SNIa observations and show that future data require the inclusion of this error in cosmological parameter estimation.
+Our method complements several existing methods that seek to estimate peculiar velocities using galaxy data and N-body simulations and extends to higher redshifts.
+Cosmology – Bayesian statistics, Peculiar motion, Type Ia supernovae.
+I Introduction
 The spatially flat $\Lambda$CDM model, based on the cosmological principle of homogeneity and isotropy, constituting cold dark matter and the cosmological constant that dominates the universe at present, has become the standard description of our observed universe.
-It has been remarkably successful in describing the observed anisotropies in the Cosmic Microwave Background (CMB)[1], large-scale clustering of matter [2, 3, 4], and the background expansion history inferred from the redshift and brightness measurements of Type Ia supernovae (SNIa) [5, 6, 7].
-In addition, a growing number of space and earth missions have ushered in the era of precision cosmology [8], which allows us to constrain cosmological parameters at a sub-percent level of precision.
-For example, the measurement of the CMB temperature and polarization anisotropies by the Planck mission has allowed us to constrain the Hubble constant within an uncertainty of $0.5$ km/s/Mpc [1].
+It has been remarkably successful in describing the observed anisotropies in the Cosmic Microwave Background (CMB)Aghanim et al.
+(2020), large-scale clustering of matter Kwan et al.
+(2017); Cole et al.
+(2005); Adame et al.
+(2024), and the background expansion history inferred from the redshift and brightness measurements of Type Ia supernovae (SNIa) Conley et al.
+(2010); Brout et al.
+(2022); Scolnic et al.
+(2022).
+In addition, a growing number of space and earth missions have ushered in the era of precision cosmology Turner (2022), which allows us to constrain cosmological parameters at a sub-percent level of precision.
+For example, the measurement of the CMB temperature and polarization anisotropies by the Planck mission has allowed us to constrain the Hubble constant within an uncertainty of $0.5$ km/s/Mpc Aghanim et al.
+(2020).
 This increasing quality and quantity of data has enabled us to extract more detailed information about the present and past of the universe.
-However, the onset of the era of greater precision has also revealed systematic discrepancies in the standard model of cosmology, such as the observed tensions in $H_{0}$ and $S_{8}$ [9, 10, 11].
-Taking these discrepancies as a hint of new physics beyond the standard $\Lambda$CDM model, several modifications to the theory have been proposed in the literature, on the contrary, untreated or systematic errors in the analysis as a source of observed discrepancies remain a less explored but plausible alternative.
-For example, there are certain anomalies in the CMB observations, such as lensing anomalies, whose presence depends on the method used for statistical analysis.
-For example, the standard Plik and Commander likelihood of Planck 2018 data prefer a higher value of the lensing parameter $A_{L}$ than the prediction of the $\Lambda\rm CDM$ model, $A_{L}=1$.
-However, in the likelihood NPIPE based on the same Planck data but different modeling of systematics, there is no such anomaly [12].
-It was shown in [13] that alleviating these tensions with extensions to the $\Lambda$CDM model is not sufficient to claim a hint for physics beyond the $\Lambda$CDM model.
+However, the onset of the era of greater precision has also revealed systematic discrepancies in the standard model of cosmology, such as the observed tensions in $H_{0}$ and $S_{8}$ Di Valentino et al.
+(2021a, b); Abdalla et al.
+(2022).
+Taking these discrepancies as a hint of new physics beyond the standard $\Lambda$CDM model, several modifications to the theory have been proposed in the literature.
+However, untreated or systematic errors in the analysis as a source of observed discrepancies remain a less explored but plausible alternative.
+For instance, there are certain anomalies in the CMB observations, such as lensing anomalies, whose presence depends on the method used for statistical analysis.
+Specifically, the standard Plik and Commander likelihood of Planck 2018 data prefer a higher value of the lensing parameter $A_{L}$ than the prediction of the $\Lambda\rm CDM$ model, $A_{L}=1$.
+However, in the likelihood NPIPE based on the same Planck data but different modeling of systematics, there is no such anomaly Rosenberg et al.
+(2022).
+It was argued in Cortês and Liddle (2024) that alleviating these tensions with extensions to the $\Lambda$CDM model is not sufficient to claim a hint for physics beyond the $\Lambda$CDM model.
 This also highlights the fact that increasing the precision in measurement also requires more accurate statistical analysis techniques to obtain consistent results from different observational probes.
@@
 This is usually justified if the errors in the subsidiary observables are small compared to the errors in the dependent variable.
-In statistics literature, the study of model fitting when there are errors in both dependent and independent variables is called errors-in-variables models [14].
+In statistics literature, the study of model fitting when there are errors in both dependent and independent variables is called errors-in-variables models Dougherty (2007).
 Many different methods have been proposed to deal with this issue if additional information can be obtained about the independent variable.
-The instrumental variable method [14] considers another variable that is related to the true regressor with some known relation.
-The method of repeated measurement [14] uses multiple observations of the same variable to minimize error.
-For the special case, when the model is a monotonic and differentiable function of the regressor variable, the likelihood can be written by inverting the model function [15].
+The instrumental variable method Dougherty (2007) considers another variable that is related to the true regressor with some known relation.
+The method of repeated measurement Dougherty (2007) uses multiple observations of the same variable to minimize error.
+For the special case, when the model is a monotonic and differentiable function of the regressor variable, the likelihood can be written by inverting the model function Gregory (2005).
 However, there is no general method for fitting arbitrary errors-in-variable models.
@@
 Their impact can be detected in the redshift-space two-point correlation function of galaxies (e.g.
-[16] and the references therein for details).
+Dodelson and Schmidt (2020) and the references therein for details).
 These velocities can also be detected statistically using the kSZ (Kinetic Sunayaev-Zeldovich) effect (e.g.
-[17, 18]).
+Aghanim et al.
+(2018); Tanimura et al.
+(2021)).
 The line-of-sight peculiar motion of host galaxies of supernovae contributes to the total observed redshift and cannot be disentangled from the contribution from cosmological expansion [^2] The contribution of peculiar motion also arises from our motion with respect to the cosmic rest frame.
-The magnitude and direction of this component can be determined from the measurement of the CMB dipole.
+The magnitude and direction of this component can be determined by measuring the CMB dipole.
 It is corrected for in the available SNIa data as in the Pantheon sample..
@@
 The impact of peculiar velocities could be more than an order of magnitude larger than the instrumental uncertainty but still small compared to Hubble recessional velocities except in the nearby universe.
-The impact of peculiar motion on the cosmological parameters has been studied in the literature using N-Body simulations [19, 20, 21, 22].
-The peculiar velocity correction in the Pantheon+ SNIa sample was obtained in [23].
-These studies focus mainly on the determination of the Hubble constant.
-In this work, we analyze the impact of peculiar velocities on the determination of cosmological parameters using SNIa data by studying the problem of fitting models with errors in both dependent and independent variables.
-We analyze the Pantheon SNIa sample [24] using three estimators—one without peculiar motions, one based on a locally linear magnitude–redshift relationship, and finally, one that accounts for the theoretical model without any approximation.
-We then apply our estimators on simulated data with parameters (e.g.
-number of SNIa, errors on magnitude) typical of the current and future SNIa data sets.
-For our study, we consider two cosmological models: a spatially flat $\Lambda$CDM model and a $w$CDM model with constant $w$.
+The impact of peculiar motions on the estimation of cosmological parameters using SNIa data has been studied in the literature using N-Body simulations (e.g.
+Gavas et al.
+(2024); Wu and Huterer (2017); Wojtak et al.
+(2014); Peterson et al.
+(2024)) and galaxy data (e.g.
+Davis et al.
+(2011); Peterson et al.
+(2022); Giani et al.
+(2024)).
+In particular, Peterson et al.
+(2022) modeled peculiar velocities within the framework of the $\Lambda$CDM model as a sum of a large-scale correlated component and a small-scale random component and estimated them independently of the SNIa data using complementary astrophysical probes and finally estimated their impact on the Hubble parameter.
+Using a frequentist technique, Davis et al.
+(2011) modeled the peculiar velocity variance as an additional term in the covariance matrix for the magnitude errors of SNIa to compute the corrections these velocities induce in the estimation of cosmological parameters such as the Hubble constant $H_{0}$ and the dark energy equation of state parameter $w$ from the SNIa data.
+This method is somewhat similar to one of our estimators based on an approximate locally linear magnitude–redshift relationship discussed in Section IV.
+However, our estimator uses the slope of the magnitude-redshift relationship to propagate redshift errors into the likelihood function, while Davis et al.
+(2011) used a constant term.
+Our main approach to accounting for peculiar velocities in the SNIa data is more direct and makes minimal assumptions about the origin of this motion.
+Furthermore, our method does not require additional input from independent astrophysical probes of peculiar motion.
+It is based on treating each velocity as an independent parameter drawn from a flat prior centered at the position of the measured redshift, which means that no specific underlying cosmological model is assumed for peculiar motion.
+Each velocity is estimated using a Bayesian method based on likelihood function and then marginalized to assess their impact on cosmological parameters.
+To this end, we develop a Bayesian method that is applicable to the problem of fitting general non-linear models with errors in both dependent and independent variables.
+We analyze the Pantheon SNIa sample Scolnic et al.
+(2018) using three estimators: 1) one without including peculiar motions, 2) one based on a locally linear magnitude–redshift relationship, with an additional assumption of Gaussian errors in redshift, and finally 3) one that accounts for peculiar motions without making any approximation.
+The second estimator was considered for comparison with the more general third estimator, as it is computationally much less expensive.
+To evaluate the efficacy of our estimators, we also applied our estimators to synthetic data with parameters (e.g., number of SNIa, errors on magnitude) typical of the current and future SNIa data sets.
+This is useful to establish the efficacy of our estimators for the current quality of data and also to demonstrate that higher-quality future data sets would lead to incorrect estimates of cosmological parameters without proper accounting of peculiar motions.
+For our study, in all cases we consider two cosmological models: a spatially flat $\Lambda$CDM model and a $w$CDM model with constant $w$.
 This paper is organized as follows.
-In Section 2, we discuss in general our method for dealing with the problem of fitting arbitrary models with errors in both variables.
-In Section 3, we present the magnitude–redshift relation as an errors-in-variables model.
-In Section 4, we apply this method to estimate cosmological parameters from supernova data for the $\Lambda$CDM and the $w$CDM model, and discuss the modification in the results as compared to the standard model fitting method that does not consider the contribution of peculiar motion.
-In Section 5, we conclude with a discussion of the results and potential future applications of our method.
+In Section II, we discuss in general our method for dealing with the problem of fitting arbitrary models with errors in both variables.
+In Section III, we present the magnitude–redshift relation as an error-in-variables model.
+In Section IV, we apply this method to estimate cosmological parameters from supernova data for the $\Lambda$CDM and the $w$CDM model, and discuss the modification in the results as compared to the standard model fitting method that does not consider the contribution of peculiar motion.
+In Section IV, we conclude with a discussion of the results and potential future applications of our method.
 In Appendix A, we derive an expression for likelihood in the linear approximation of the magnitude–redshift relation.
-2 Methodology
-The problem of regression with an arbitrary errors-in-variables model is interesting and, to the best of our knowledge, not a commonly addressed problem in astronomy.
+II Methodology
+The problem of regression with an arbitrary error-in-variables model is an interesting problem and, to the best of our knowledge, not a commonly addressed problem in astronomy.
 Since astronomical data usually have errors in independent variables, this could have general applications in a variety of problems in astronomy.
-In this section, we describe a Bayesian approach to this problem.
-In Section 2.1, we describe our method in general while motivating our choice of the particular form of estimator that we propose.
-Then, in Section 3, we specialize it for the estimation of cosmological parameters from the magnitude–redshift data of the Pantheon sample of supernovae [24].
+In this section, we describe a Bayesian solution to this problem.
+In Section II.1, we begin by describing the general background of the problem and develop our estimator by motivating it with the familiar case of independent variables without errors.
+Then, in Section III, we specialize our estimator for estimating cosmological parameters from the magnitude–redshift data of the Pantheon sample of supernovae Scolnic et al.
+(2018).
 We consider the contribution of peculiar motion of supernova host galaxies to the redshift uncertainty, which leads to its formulation as an errors-in-variables problem.
-We also compare it to an approach that uses a local linear approximation of the magnitude–redshift relation and uses known methods of handling errors-in-variables for linear models.
-2.1 Estimator for Errors-in-Variables
-Quite often, the output of experiments is a pair $(x_{i},y_{i})$ [^3] It is worth noting that in general, the data could have several independent variables $(x_{i},y_{i},z_{i},\cdots)$.
-It is straightforward to generalize this analysis to such multivariate cases.
-of $N$ data points with uncertainties, assuming the randomness to be Gaussian, $\sigma_{x_{i}}$ and $\sigma_{y_{i}}$, associated with each observed value $x_{i}$ and $y_{i}$ respectively.
-The quantity $y_{i}$ often depends on the independently measured quantity $x_{i}$ through a model that could be given through a mathematical relation
-$y_{i}=f(\theta,x_{i}),$ (2.1)
-or, in general, could be given by an algorithm such as a numerical simulation or a complex mathematical operation.
-Here, $f(\theta,x_{i})$ is a function of the $m$ model parameters, $\theta\equiv(\theta_{1},\theta_{2},...\theta_{m})$ and the variable $x_{i}$.
-In our notation, if any expression depends explicitly on a single variable, then we notate it with a subscript; else, we use unscripted symbols to refer to them collectively.
-We shall soon focus on the case of apparent magnitude versus redshift data that comprises the measured pair $(z_{i},m_{i})$, which might be helpful to the reader while reading the following discussion to make it less abstract.
-If the function $f(\theta,x_{j})$ is a linear function of $x_{j}$, then estimating the parameters $\theta$ from data that have errors on both axes is quite straightforward (see, for example, [25] and the derivation in Appendix A).
+Later in this paper, we have also compared it to an approximate method that uses a local linear approximation of the magnitude–redshift relation and assumes a Gaussian distribution of errors due to peculiar motions.
+The development of this method can be found in the Appendix A.
+This comparison is important because the approximate method is computationally inexpensive and is therefore useful for very large data sets.
+Our results show that, despite its assumptions, it gives comparable results to the exact method.
+However, the exact estimator is more flexible in handling a greater variety of problems.
+II.1 Estimator for Errors-in-Variables
+Frequently, experiments produce $N$ pairs $(x_{i},y_{i})$ of noisy data points with uncertainties (assuming the randomness to be Gaussian) $\sigma_{x_{i}}$ and $\sigma_{y_{i}}$ associated with each observed value $x_{i}$ and $y_{i}$, respectively.
+The quantity $y_{i}$ often depends on the independently measured quantity $x_{i}$ through a physical model that is often given in terms of a mathematical relation,
+$y_{i}=f(\theta,x_{i}),$ (1)
+or, in general, could be given by an algorithm such as a numerical code or a multistep mathematical operation that outputs $y$ for every $x$.
+However, our treatment is based on expressing the association of each $x$ with $y$ abstractly through the above relationship.
+The function $f(\theta,x_{i})$ is generally a function of $m$ unknown model parameters $\theta\equiv(\theta_{1},\theta_{2},...\theta_{m})$ along with the variable $x_{i}$.
+In the notation used in this paper, if any expression depends explicitly on a single variable, then we notate it with a subscript; else, we use unscripted symbols to refer to them collectively.
+We shall soon focus on the case of apparent magnitude-redshift data that comprise the measured pair $(z_{i},m_{i})$.
+It might be helpful to the reader to keep this concrete model in mind while reading this section.
+The parameters $\theta$ would then be the usual cosmological parameters such as the Hubble constant, various matter energy densities, and the equation of state parameter.
+The most straightforward case is where the independent variable $x$ can be measured with negligible error.
+This is often how such data are fitted to models, often ignoring the $x$ errors even if they are not small.
+This is due to the fact that the canonical method relies on working with the likelihood function derived from the errors in the dependent variable $y$.
+Therefore, we must find a way to incorporate $x$ errors as an additional source of error in the likelihood function.
+If the function $f(\theta,x_{j})$ is a linear function [^3] Although in general a nonlinear model could refer to a model $f(\theta,x_{j})$ with nonlinear dependence on $\theta$ or $x_{i}$, in this paper by a nonlinear model we always refer to the latter case, of $x_{j}$, then estimating the parameters $\theta$ from data that have errors in both variables is well established and quite straightforward (see, for example, Press et al.
+(1992) and the derivation in Appendix A).
 This method is based on the conversion of the $x$ error into the $y$ error through the slope of the linear function and adding it in quadrature to the $y$ error while constructing the likelihood function.
 However, when the model is nonlinear in $x_{j}$, this trick works only by approximating the fitting function as locally a linear function, as shown for the special case of the magnitude–redshift relation in the Appendix A.
-Although, in general, a nonlinear model could refer to nonlinear dependence on $\theta$ or $x_{i}$, in this paper, by nonlinear model, we always refer to the latter case.
-2.1.1 Estimator for nonlinear models
-To keep things general, let $P_{\rm Y}(\epsilon_{Y})$ be the probability distribution function for random statistical noise in the measurement of $y$, so that the observed values are offset from their true value by a random number, $y_{i}=f(\theta,x_{i})+\epsilon_{Yi}$, where $\epsilon_{Yi}$ is the random error in the measurement $y_{i}$.
-Therefore, the likelihood function is given by
-$\mathcal{L}(\mathcal{D}\mid\theta)=\prod_{i}P_{Y}(y_{i}-f(\theta,x_{i})),$ (2.2)
+Our aim in this paper is to develop a method that is not approximate or ad hoc and treats errors in $x$ and $y$ on equal footing.
+This turns out to be possible in Bayesian statistics, as we show next.
+II.1.1 Estimator for nonlinear models
+We introduce our approach through the familiar case of data with errors only in the dependent variable $y$ and none in the independent variable $x$.
+Let the true values of $y$ be denoted by $y^{*}$.
+The observed values of $y$ differ from their true value by a random number, $y_{i}=y^{*}_{i}+\epsilon_{Yi}$, where $\epsilon_{Yi}$ is the random error in measurement $y_{i}$.
+Since $x$ are measured without error, $y^{*}_{i}=f(\theta,x_{i})$.
+If $P_{\rm Y}(\epsilon_{Y})$ is the probability distribution function for random noise in the measurement of $y$, the likelihood function is given by
+$\mathcal{L}_{Y}(\mathcal{D}\mid\theta,y^{*})=\prod_{i}P_{Y}(y_{i}-y^{*}_{i})\delta(y^{*}_{i}-f(\theta,x_{i})),$ (2)
 where $\mathcal{D}$ collectively refers to the pair $(x,y)$.
-In Bayesian statistics, using Bayes’ theorem, this likelihood function can be used to write the posterior probability distribution for parameters $\theta$ as
-$P(\theta\mid\mathcal{D};I)\propto\mathcal{L}(\mathcal{D}\mid\theta;I)P(\theta\mid I)$ (2.3)
-where $P(\theta\mid I)$ is the prior probability distribution function for $\theta$, and the addition of $I$ in various terms symbolizes all the prior information that we may have (for details, see, e.g., [26]).
-Until now, the variables $x$ have been considered given and fixed.
-To incorporate error in variables $x$ in the posterior distribution, we now turn them into undetermined parameters $x^{*}$, equivalent to parameters $\theta$, in terms of which the likelihood function can be written as
-$\mathcal{L}(\mathcal{D}\mid\theta,x^{*};I)=\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))$ (2.4)
-Note that conversion of $x$ to parameter means that now the data $\mathcal{D}$ refers only to $y$.
-The information about errors in $x$ and their measured values can be brought into our analysis as a prior on $x^{*}$ through [^4] We can derive the same expressions by noting that the joint probability of errors $\epsilon_{X}$ and $\epsilon_{Y}$ is given by $P_{X}(\epsilon_{X})P_{Y}(\epsilon_{Y})$, thus the errors on $x$ can be considered to be a part of likelihood of a data set comprising the pair $(x,y)$
-$P(x^{*}_{i}\mid I)=P_{\rm X}(x^{*}_{i}-x_{i}),$
-where $P_{\rm X}$ is the probability distribution for errors in variable $x$ (assumed for simplicity to be the same for all data points, but this can be easily generalized).
-The modified posterior distribution is given by
-$P(\theta,x^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))P(x^{*}_{i}\mid I)P(\theta\mid I)$ (2.5)
-If the random noise is Gaussian, then the log of the likelihood function is given by
-$\ln\mathcal{L}(\mathcal{D}\mid\theta,x^{*})=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{y_{i}-f(\theta,x^{*}_{i})}{\sigma_{y_{i}}}\right]^{2}\equiv-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2},$ (2.6)
-and the log of the prior probability distribution function for $x$ is given by
-$\ln{P(x^{*}_{i}\mid I)}=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{x^{*}_{i}-x_{i}}{\sigma_{x_{i}}}\right]^{2}\equiv-\frac{\chi^{2}_{X}(x^{*})}{2}$ (2.7)
-In terms of these variables, the posterior distribution can be written as
-$P(\theta,x^{*}\mid\mathcal{D})\propto\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)$ (2.8)
-We can marginalize $x^{*}$ to obtain the probability distribution for $\theta$
-$P(\theta\mid\mathcal{D})\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)$ (2.9)
-In practice, marginalization is more conveniently carried out through MCMC, as we discuss later.
-Since the total number of parameters now exceeds the number of data points, there is, in principle, the possibility of overfitting the data.
-In order to obviate this possibility, we add a regularization term as a prior probability that our physical model should be such that the mean value of $\chi_{Y}^{2}$ is close to degrees of freedom of the model [^5] This expectation is technically true only for linear models, but holds quite well in general for small errors as models can then be roughly treated as locally linear..
-For the case of Gaussian noise in the dependent variable, this is just the probability distribution function for $\chi_{Y}^{2}$ given by
-$P_{\chi^{2}}(\chi^{2}_{Y})=\frac{\chi_{Y}^{k-2}\exp(-\chi^{2}_{Y}/2)}{2^{k/2}\Gamma(k/2)},$ (2.10)
-where $k=N-m$ is the desired degrees of freedom of the fit (not accounting for noise in the $x$ values).
-For example, for the case of fitting luminosity distance to distant supernovae, this would be the total number of supernovae minus the number of cosmological parameters being fitted.
+In this form, the likelihood function can only be evaluated if we know the values of $\theta$ and $y^{*}$.
+The Dirac delta function ensures that the likelihood is non-zero only when $y^{*}_{i}=f(\theta,x_{i})$.
+Since $x_{i}$ is known but not $y^{*}_{i}$, this equation can be satisfied for any value of $y^{*}_{i}$ at $\theta$ differing from their true value.
+Furthermore, the subscript $Y$ in the likelihood function has been added to make it explicit that it is derived from the random noise in $y$.
+In a sense, we can think of the true values $y^{*}$ as additional parameters of the model, which is how they are treated in the following analysis.
+Using Bayes’ theorem, this likelihood function can be used to write the posterior probability distribution for the parameters $\theta$, and $y^{*}$ as
+$P(\theta,y^{*}\mid\mathcal{D};I)\propto\mathcal{L}_{Y}(\mathcal{D}\mid\theta,y^{*})P(\theta\mid I)$ (3)
+where $P(\theta\mid I)$ is the prior probability distribution function for $\theta$, and the addition of $I$ in various terms symbolizes all the prior information that we may have (for details, see, e.g., Sivia (1996)).
+We also assumed that there is no prior information on the true values $y^{*}$.
+Combining Eq.
+(2) with this equation we obtain
+$P(\theta,y^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-y^{*}_{i})\delta(y^{*}_{i}-f(\theta,x_{i}))P(\theta\mid I)\,.$ (4)
+This posterior probability distribution has information about the true values $y^{*}$, which is of little use.
+Usually, the parameters $\theta$ are of interest; therefore, we can marginalize over $y^{*}$ by integrating over them to obtain the most commonly used expression for the posterior probability for a model with errors only in dependent variable,
+$P(\theta\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-f(\theta,x_{i}))P(\theta\mid I)\,.$ (5)
+To generalize this to the case with errors in both variables, in analogy with $y$ we write $x_{i}=x^{*}_{i}+\epsilon_{Xi}$, where $x^{*}$ is the unknown true value of $x$ in terms of which the likelihood function can be written as
+$\mathcal{L}_{XY}(\mathcal{D}\mid\theta,x^{*},y^{*})=\prod_{i}P_{Y}(y_{i}-y^{*}_{i})P_{X}(x_{i}-x^{*}_{i})\delta(y^{*}_{i}-f(\theta,x^{*}_{i})),$ (6)
+where we have used the fact that $x$ and $y$ are statistically independent, therefore, their joint probability can be obtained simply by multiplying their individual probability distribution functions.
+However, note that since $x$ are now noisy, the argument of delta function now contains only true values $(x^{*},y^{*})$.
+The modified posterior distribution is now given by
+$P(\theta,x^{*},y^{*}\mid\mathcal{D};I)\propto\prod_{i}P_{Y}(y_{i}-y^{*}_{i})P_{X}(x_{i}-x^{*}_{i})\delta(h(x^{*},y^{*};\theta))P(\theta\mid I)$ (7)
+where we have defined $h(x^{*},y^{*};\theta)\equiv y^{*}_{i}-f(\theta,x^{*}_{i})=0$ as a symmetric way to write the model relationship between $x$ and $y$.
+The obvious advantage is that it makes the difference between dependent and independent variables disappear.
+It is worth noting that in general, the data could have several independent variables $(x_{i},y_{i},z_{i},\cdots)$ related through $h(x_{i},y_{i},z_{i},\cdots)=0$.
+It is straightforward to generalize this analysis to such multivariate cases.
+The posterior distribution in Eq.
+(7) is a function not only of the model parameters but also of the unknown true values of $x$ and $y$ for each data point.
+If we are interested only in the model parameters $\theta$, then we can integrate over additional parameters.
+This yields
+$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\prod_{i}P_{Y}(y_{i}-f(\theta,x^{*}_{i}))P_{X}(x_{i}-x^{*}_{i})P(\theta\mid I)$ (8)
+If we assume Gaussian errors in both variables with standard deviations $\sigma_{y}$ and $\sigma_{x}$, respectively, then after defining
+$\chi^{2}_{Y}(\theta,x^{*})=\sum_{i=1}^{N}\left[\frac{y_{i}-f(\theta,x^{*}_{i})}{\sigma_{y_{i}}}\right]^{2},\quad\chi^{2}_{X}(x^{*})=\sum_{i=1}^{N}\left[\frac{x^{*}_{i}-x_{i}}{\sigma_{x_{i}}}\right]^{2}\,,$ (9)
+the marginalized posterior distribution can be written as
+$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\theta,x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\theta\mid I)\,.$ (10)
+This form is used in Appendix A to derive an approximate form of the estimator where integration over $x^{*}$ can be done analytically.
+However, if we do not assume a specific form for $P_{X}$ then assuming Gaussian errors only in $y$ we get
+$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2}\right]P_{X}(x_{i}-x^{*}_{i})P(\theta\mid I)\,.$ (11)
+In practice, the integration in Eq.
+(11) is more conveniently carried out with Markov Chain Monte Carlo (MCMC), as we discuss later.
+Since the total number of model parameters now exceeds the number of data points, there is, in principle, the possibility of overfitting the data.
+In order to obviate this possibility, we add a regularization term to ensure that the physical model satisfy the condition that the mean value of $\chi_{Y}^{2}$ is close to the degrees of freedom of the model [^4] This expectation is technically true only for linear models, but holds well in general for small errors as models can then be roughly thought of as locally linear..
+The marginalized $x^{*}$ can be considered as subsidiary parameters not directly related to the physical model.
+Therefore, the degrees of freedom is given by the number of data points, $N$, minus the number of physical parameters, $m$.
+For the case of Gaussian noise in the dependent variable, the required regularization term is just the probability distribution function for $\chi_{Y}^{2}$ given by
+$P_{\chi^{2}}(\chi^{2}_{Y})=\frac{\chi_{Y}^{k-2}\exp(-\chi^{2}_{Y}/2)}{2^{k/2}\Gamma(k/2)},$ (12)
+where $k=N-m$ is the desired degrees of freedom of the fit.
+For the case of fitting magnitude-redshift relationship for distant supernovae, this would be the total number of supernovae minus the number of cosmological parameters being fitted.
 Therefore, our final form for the posterior distribution function for $\theta$ can be obtained by multiplying it with the probability distribution function for $\chi_{Y}^{2}$
-$P(\theta\mid\mathcal{D})\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\bm{\theta},x^{*})+\chi^{2}_{X}(x^{*})}{2}\right]P(\bm{\theta}\mid I)P_{\chi^{2}}(\chi_{Y}^{2})\,.$ (2.11)
-Regularization is commonly used in fitting complex models with a large number of parameters as it acts as a constraint and facilitates stable maximization of the posterior distribution (e.g.
-[26]).
-However, as we note later, the addition or omission of this term may not always change the dimensionality of the fit, as explained further in the context of the specific cosmological application of this method presented in the paper.
-3 Magnitude–Redshift relation
-Although the measurement error in the observed redshift of a supernova host galaxy is negligibly small, due to the possible peculiar motion of the host galaxy with respect to an observer at rest with respect to the universe, their observed redshift is shifted with respect to the theoretical cosmological redshift that encodes the expansion history of the universe.
+$P(\theta\mid\mathcal{D};I)\propto\int d^{N}x^{*}\exp\left[-\frac{\chi^{2}_{Y}(\theta,x^{*})}{2}\right]P_{X}(x_{i}-x^{*}_{i})P({\theta}\mid I)P_{\chi^{2}}(\chi_{Y}^{2})\,.$ (13)
+Regularization is commonly used to fit complex models with a large number of parameters as it acts as a constraint and facilitates stable maximization of the posterior distribution (e.g.
+Sivia (1996)).
+However, as we note later, the addition or omission of this term may not always change the dimensionality of the fit, as is further explained in the context of the specific cosmological application of this method presented in the paper.
+It is clear that the method advocated here is more expensive to implement numerically since each independent variable is a parameter.
+We chose cosmological data from SNIa as a proof of concept for our proposed method.
+III Magnitude–Redshift Relation
+Although the measurement error in the observed redshift of a supernova host galaxy is negligibly small, due to the possible peculiar motion of the host galaxy relative to a comoving observer, its observed redshift is shifted with respect to the theoretical cosmological redshift that encodes the expansion history of the universe.
 To estimate this shift, we note that
-$\frac{\nu_{0}}{\nu_{\rm obs}}=\frac{\nu_{0}}{\nu_{\rm com}}\frac{\nu_{\rm com}}{\nu_{\rm obs}}\approx(1+z)(1+{v_{p}}/{c})\approx 1+z+{v_{p}}/{c}$ (3.1)
-where $\nu_{\rm obs},\nu_{0}$, and $\nu_{\rm com}$ are the frequencies observed at earth, in the rest frame of the galaxy and a comoving observer at the position of the galaxy, and $v_{p}$ is the component of velocity of a galaxy with respect to the comoving observer parallel to the line-of-sight from the observer to the supernova host galaxy.
-The approximation in the last equation assumes $v_{p}/c\ll 1$, so we can neglect the transverse Doppler effect.
-For small redshifts where this correction dominates, it suffices to ignore the $zv_{p}/c$ term, leading to the final form.
+$\frac{\nu_{0}}{\nu_{\rm obs}}=\frac{\nu_{0}}{\nu_{\rm com}}\frac{\nu_{\rm com}}{\nu_{\rm obs}}\approx(1+{v_{p}}/{c})(1+z)$ (14)
+where $\nu_{\rm obs},\nu_{0}$, and $\nu_{\rm com}$ are the supernova photon frequencies observed on earth, in the rest frame of the galaxy, and by a comoving observer at the position of the galaxy; and $v_{p}$ is the component of velocity of a galaxy with respect to the comoving observer, parallel to the line of sight from the observer to the supernova host galaxy.
 Peculiar motions also cause beaming of light towards the observer, alter the observed flux, and cause additional k-correction (e.g.
-see [27] for the combined impact of all these effects on the number count).
+see Ellis and Baldwin (1984) for the combined impact of all these effects on the number count).
 The first two effects act in concert, as an object moving towards an observer would appear closer and less redshifted.
@@
 At low redshifts, this effect is enhanced as peculiar redshift/blueshift can become comparable to the redshift of the host galaxy.
-As more and more supernovae are observed, the effective statistical error in luminosity distance reduces at any redshift due to statistical averaging.
+As more and more supernovae are observed, the effective statistical error in luminosity distance reduces at any redshift because of statistical averaging.
 Thus, we need techniques that allow us to handle errors on redshift, even if they are small.
@@
 The apparent magnitude $m$ and the luminosity distance $D_{\rm L}$ of a supernova of known absolute magnitude $M$ are related through
-$m=M+5\log\left(\frac{D_{\rm L}}{\text{Mpc}}\right)+25\;.$ (3.2)
+$m=M+5\log\left(\frac{D_{\rm L}}{\text{Mpc}}\right)+25\;.$ (15)
 For a flat Friedmann–Lemaître–Robertson–Walker universe, the luminosity distance of a source at redshift $z$ can be expressed as
-$D_{\rm L}=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{h(z^{\prime})},$ (3.3)
+$D_{\rm L}=\frac{c(1+z)}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{h(z^{\prime})},$ (16)
 where the dimensionless Hubble parameter is given by
-$h(z)=\left[\Omega_{\rm M}(1+z)^{3}+\Omega_{\rm DE}(1+z)^{3(1+w)}\right]^{1/2}\,,$ (3.4)
-where $\Omega_{\rm M}$ and $\Omega_{\rm DE}$ are the present-day matter and dark energy densities respectively.
+$h(z)=\left[\Omega_{\rm M}(1+z)^{3}+\Omega_{\rm DE}(1+z)^{3(1+w)}\right]^{1/2}\,,$ (17)
+where $\Omega_{\rm M}$ and $\Omega_{\rm DE}$ are the present-day matter and dark energy densities, respectively.
 The dark energy and pressure are related through the parameter $w$ through $p_{\rm DE}=w\rho_{\rm DE}$.
@@
 The apparent magnitude can be written as
-$m=M+5\log(cH^{-1}_{0}/{\rm Mpc})+5\log\left[(1+z)\int_{0}^{z}\frac{dz^{\prime}}{h(z)}\right]+25\,.$ (3.5)
+$m=M$ $+$ $5\log(cH^{-1}_{0}/{\rm Mpc})$ (18)
+$+$ $5\log\left[(1+z)\int_{0}^{z}\frac{dz^{\prime}}{h(z)}\right]+25\,.$
 Here, $w$, $\Omega_{\rm M}$, and $H_{0}$ are free parameters of the model. However, note that $H_{0}$ is degenerate with $M$, which allows us to estimate only their combination.
-4 Results
+IV Results
 In cosmological studies, our primary concern is with the parameters $w$ and $\Omega_{\rm M}$.
-For presenting our results, we need to marginalize various parameters, which means integrating them in the posterior distribution.
+Although distant SNIa are not ideal for determining $H_{0}$ due to degeneracy with $M$, we also present results for the Hubble parameter for the $\Lambda$CDM model.
+To break the degeneracy, we use the absolute magnitude $M=-19.24$ corresponding to the SH0ES Riess et al.
+(2022) value of $H_{0}=73.04$ km/s/Mpc for the Pantheon sample.
+In our analysis, we consider three distinct estimators to present our results.
+  • The first estimator labeled $\mathcal{E}_{1}$ is based on the standard method that does not account for redshift errors and treats them as perfectly measured.
+  • The second estimator labeled $\mathcal{E}_{2}$ is based on the approximate linearized magnitude–redshift relation that includes the redshift error using the likelihood function (Eq.
+(24)) derived in the Appendix A.
+This estimator assumes that the redshift errors are Gaussian.
+  • The MCMC-based final estimator $\mathcal{E}_{3}$ uses Eq.
+(13) described in Section II.1.1, which includes the contribution of peculiar velocity to the observed redshift by treating each redshift as a uniformly distributed random number in a range around the measured value of the redshift.
+Note that only the line-of-sight component of the peculiar velocity contributes to the change in redshift.
+By the isotropy of the peculiar velocity dispersion, we have $\sigma_{v_{\rm 3D}}=\sqrt{3}\sigma_{v_{p}}$.
+Thus, any assumption we make about the magnitude of peculiar velocity must keep this in mind.
+For estimator $\mathcal{E}_{3}$, the distribution $P_{X}$ is minimally assumed to be uniform in a specified redshift range.
+More specifically, we assume that each observed $z_{i}$ in the data deviates from its true value $z^{*}_{i}$, which falls within a specified uncertainty range $z^{*}_{i}\in[z_{i}^{\rm min},z_{i}^{\rm max}]$, where the min and max values are chosen to correspond to our assumptions about peculiar velocities.
+We also need to assign a value to $\sigma_{z_{i}}$ in Eq.
+(24).
+We do that by setting the range chosen for $\mathcal{E}_{3}$ to be three times $\sigma_{z_{i}}$.
+This ensures that the one-sigma range from $-\sigma_{z_{i}}$ to $\sigma_{z_{i}}$ for a Gaussian distribution containing 99.7% probability is equal to the redshift range allowed in $\mathcal{E}_{3}$.
+Therefore, $\sigma_{z_{i}}=(z_{i}^{\rm max}-z_{i}^{\rm min})/6$
+For presenting our results, we need to marginalize over various parameters, which means integrating over them in the posterior distribution.
 Also note that in Eq.
-(2.11), we need to integrate over $x^{*}$, which corresponds to the redshift $z$ in this case.
-Clearly, it is not feasible to integrate over so many variables.
-Therefore, to carry out marginalization, we employ the Markov Chain Monte Carlo (MCMC) sampling technique.
-Specifically, we develop and implement the Metropolis-Hastings algorithm within the MCMC framework to draw samples from the posterior distribution and use GetDist [28] python library for the analysis of MCMC samples.
+(13), we need to integrate over the variable $x^{*}$, which corresponds to the redshift $z$ in this case.
+Clearly, it is not feasible to numerically integrate so many variables.
+Therefore, to carry out marginalization, we employ the MCMC sampling technique.
+Specifically, we develop and implement the Metropolis-Hastings algorithm within the MCMC framework to draw samples from the posterior distribution and use GetDist Lewis (2019) python library for the analysis of MCMC samples.
+We found that MCMC sampling requires a longer time to converge if we treat cosmological parameters and redshift parameters on par.
+However, the efficiency of convergence is significantly improved by drawing multiple samples of redshifts for the same cosmological parameters.
+Thus, in all the results below, MCMC samples the redshift five times more often than the cosmological parameters.
+This ensures that MCMC is able to optimally sample the parameter space.
 Figure 1: The figure displays one-dimensional posterior distributions and two-dimensional (1- and 2-$\sigma$) contours of $\Lambda$CDM parameters from Pantheon data using three different estimators.
-Estimator $\mathcal{E}_{1}$ (red) ignores any contribution to redshift from the peculiar motion of supernovae host galaxies, $\mathcal{E}_{2}$ considers it for the linearized model (see Appendix A for details), while $\mathcal{E}_{3}$ (green) represents our estimator, which incorporates the contribution from peculiar motion for the exact model.
-Parameters Priors Best-fit values
+Estimator $\mathcal{E}_{1}$ (red) ignores any contribution to redshift from the peculiar motion of supernovae host galaxies, $\mathcal{E}_{2}$ (blue) considers it for the linearized model (see Appendix A for details), while $\mathcal{E}_{3}$ (green) represents our estimator, which incorporates the contribution from peculiar motion for the exact model.
+The errors on the estimated parameters are larger for estimators $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$.
 […]
-$\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.286\pm 0.012$ $0.296\pm 0.014$ $0.302\pm 0.016$
-$H_{0}$ $\mathcal{U}(60,80)$ $73.59\pm 0.229$ $73.34\pm 0.287$ $73.13\pm 0.343$
-[…]
-Although distant SNIa are not ideal for determining $H_{0}$ due to degeneracy with $M$, in the following sections, we also present results for the Hubble parameter for the $\Lambda$CDM model.
-In order to break the aforementioned degeneracy, we use the absolute magnitude $M=-19.2478$ corresponding to the SH0ES [29] value of $H_{0}=73.04$ km/s/Mpc for the Pantheon sample.
-For the redshift error, we assume that each observed $z_{i}$ in the data deviates from its true value $z^{*}_{i}$, which falls within a specified uncertainty range $z^{*}_{i}\in[z_{i}^{\rm min},z_{i}^{\rm max}]$, where the min and max values are chosen to correspond to our assumptions about peculiar velocities.
-This is imposed even if we assume the distribution of redshift to be nonuniform.
-From analytical theory or numerical simulations, we can estimate the variance of galaxy peculiar velocities as a function of redshift.
-Also, note that only the line-of-sight component of the peculiar velocity contributes to the change in redshift.
-By the isotropy of the peculiar velocity dispersion, we have $\sigma_{v_{\rm 3D}}=\sqrt{3}\sigma_{v_{p}}$
-In our analysis below, we consider three distinct cases to present our results.
-The standard case is labeled as $\mathcal{E}_{1}$, where we do not account for the redshift error.
-In Appendix A, we derive an estimator based on the linearized magnitude–redshift relation labeled $\mathcal{E}_{2}$ to include the redshift error using the likelihood function (Eq.
-(A.6)).
-This is of interest since the MCMC-based final estimator ($\mathcal{E}_{3}$) using the method described in Section 2.1.1 (Eqs.
-(2.2)–(2.11)), which includes the contribution of peculiar velocity to the observed redshift by treating each redshift as a parameter, can be computationally expensive for large data sets.
-4.1 Application to Pantheon Data
-We apply our method to fit the $\Lambda$CDM and the $w$CDM models to the magnitude–redshift data of Type Ia supernovae in the Pantheon sample [24].
+${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
+$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.285\pm 0.013$ $0.296\pm 0.015$ $0.306\pm 0.016$
+$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.86\pm 0.23$ $73.59\pm 0.30$ $73.36\pm 0.31$
+Table 1: Priors and best-fit values for the spatially-flat $\Lambda$CDM model using Pantheon data.
+Results are shown for three different estimators: $\mathcal{E}_{1}$ ignores peculiar velocities, $\mathcal{E}_{2}$ includes them in the linear approximation, and $\mathcal{E}_{3}$ incorporates them exactly.
+IV.1 Application to Pantheon Data
+We apply our method to fit the $\Lambda$CDM and the $w$CDM models to the magnitude–redshift data of SNIa in the Pantheon sample Scolnic et al.
+(2018).
 This data set contains 1048 spectroscopically confirmed SNIa.
-For all cases, we run 12 independent MCMC chains for $10^{6}$ steps.
+Using $\Delta z=({v_{p}/c})(1+z)$, the standard deviation used in $\mathcal{E}_{2}$ is given by $\sigma_{z}=(\sigma_{v_{p}}/c)(1+z)$.
+Throughout, we use $\sigma_{v_{p}}=1000\;\mathrm{km/s}$.
+For the estimator $\mathcal{E}_{3}$, a uniform prior $[-3000,3000]\;\rm km/s$ has been used for the peculiar velocity $v_{p}$.
+For all cases, we run 12 independent MCMC chains for $10^{5}$ steps.
 We use the Gelman-Rubin criterion of convergence.
-For the $\Lambda$CDM model, we achieve $R-1<0.001$, while for the $w$CDM model it is $R-1<0.005$.
-For $z$-errors, we use uniform priors for $v_{pi}$ in the range: $-1000\leq v_{pi}\leq 1000\,\rm km\,s^{-1}$ around the measured redshift.
-To verify convergence and test its rate, we increase the range of peculiar velocities to $-3000\leq v_{pi}\leq 3000\,\rm km\,s^{-1}$.
-Increasing this interval reduces the acceptance rate and increases $R-1$ for the fixed number of MCMC steps.
-However, $R-1$ is always less than $0.01$.
-Thus, it leads to slower convergence, but our results remain unchanged; we also obtain the same results by assuming a Gaussian distribution for peculiar velocities, imposing only the condition that $x^{*}>0$.
-Figure 2: The figure shows posterior distributions and contours of $w$CDM parameters from Pantheon data using three different estimators. It follows the convention used in Figure 1.
-Parameters Priors Best-fit values
-[…]
-$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.6)$ $0.348\pm 0.035$ $0.335\pm 0.043$ $0.330\pm 0.046$
-$H_{0}$ $\mathcal{U}(60,80)$ $74.07\pm 0.368$ $73.80\pm 0.546$ $73.58\pm 0.703$
-$w$ $\mathcal{U}(-2,0)$ $-1.24\pm 0.14$ $-1.17\pm 0.17$ $-1.14\pm 0.18$
-[…]
-It should be noted that our results are insensitive to regularization (Eq.
-(2.11)).
-The regularization term only changes the acceptance rate, indicating that the inclusion of redshift errors does not increase the dimensionality of our fitted physical model.
-This is not surprising if we note that overfitting requires the fitting function to be able to pass through most of the noisy data points.
+For estimators $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, we achieve convergence with $R-1<0.01$, while for $\mathcal{E}_{3}$ we obtain $R-1<0.1$.
+It might be surprising that we are able to find the optimal solution even with more than a thousand parameters.
+To achieve this, we need to start our MCMC runs from the observed values of redshifts.
+However, as long as the redshifts are assumed to have peculiar velocity components not exceeding 3000 km/s, the starting point does not matter too much.
+The extension of this interval reduces the acceptance rate of MCMC and increases $R-1$ for a fixed number of steps.
+Thus, it leads to slower convergence.
+Since the cosmological line-of-sight peculiar velocities are expected to be much smaller than 3000 km/s, we have not investigated convergence by relaxing our assumptions.
+Moreover, as noted before, for faster convergence, we split the parameter space into cosmological and redshift spaces and sample the redshift space five times more frequently than the cosmological parameter space.
+We find that our results are not altered significantly by regularization (Eq.
+(13)).
+The regularization term only changes the acceptance rate and marginally changes the final $\chi^{2}$ per degree of freedom, indicating that the inclusion of redshift errors does not increase the dimensionality of our fitted physical model.
+This is not surprising if we note that overfitting requires the fitting function to be able to pass through most of the noisy data points that are displaced vertically due to errors in magnitude.
 This would require the fitting function to have the flexibility that Eq.
-(3.2) does not have, even if we convert the redshift to a variable.
-The function remains a monotonic function of redshift and thus does not cause overfitting of the data.
+(15) does not have with respect to treating the redshift as a parameter.
+The function remains monotonic, ensuring that the convergence is robust even without regularization.
 Figures 1 and 2 display the posterior distribution and the confidence regions $1\sigma$ and $2\sigma$ for the $\Lambda$CDM and $w$CDM models for the Pantheon data.
-Different colors represent the three different estimators used.
-The red color represents the standard case, $\mathcal{E}_{1}$, blue represents the linearized estimator, $\mathcal{E}_{2}$, and the green represents the MCMC estimator, $\mathcal{E}_{3}$.
+Estimators are denoted by different colors.
+The red color represents the standard case, $\mathcal{E}_{1}$, blue represents the linear approximation estimator, $\mathcal{E}_{2}$, and green represents the MCMC estimator, $\mathcal{E}_{3}$.
 In both cosmological models, we have assumed uniform priors for the cosmological parameters, which are listed in Tables 1 and 2.
-Perusal of Figures 1 & 2 indicates that accounting for peculiar velocities in the current data does not significantly change the estimated parameters.
-However, note that in Table 2 for the $w$CDM model, the marginalized estimate for the dark energy parameter $w$ is found to be marginally more consistent with the $\Lambda$CDM model, indicating that even at the current level of precision, the data could indicate non-trivial dark energy if systematic errors in redshift are not corrected.
+Perusal of Figures 1 and 2 indicates that accounting for peculiar velocities in the current data does not significantly change the estimated parameters.
+Its main impact is a small increase in errors in the estimated parameters.
+We further note that in Table 2 for the $w$CDM model, the marginalized estimate for the dark energy parameter $w$ is found to be marginally more consistent with the $\Lambda$CDM model, indicating that even at the current level of precision, the data could indicate evolving dark energy if systematic errors in redshift are not corrected.
 As the number of supernovae increases, this correction is likely to become even more significant.
 Next, we consider simulated data to investigate this issue further.
-Figure 3: Constraints on the $\Lambda$CDM parameters from synthetic data with magnitude error $\Delta m=0.2$ (left) and $\Delta m=0.05$ (right) The colour scheme of Figure 1 is followed.
-Models Parameters Priors True values Best-fit values
+Figure 2: The figure shows posterior distributions and contours of $w$CDM parameters from Pantheon data using three different estimators. It follows the convention used in Figure 1.
 […]
-$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.217\pm 0.014$ $0.240\pm 0.016$ $0.280\pm 0.020$
-$H_{0}$ $\mathcal{U}(60,80)$ $73.0$ $75.60\pm 0.338$ $74.88\pm 0.391$ $73.45\pm 0.482$
-$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.368\pm 0.023$ $0.356\pm 0.028$ $0.320\pm 0.048$
-$H_{0}$ $\mathcal{U}(60,80)$ $73.0$ $78.01\pm 0.642$ $77.09\pm 0.849$ $74.29\pm 1.024$
-$w$ $\mathcal{U}(-3,0)$ $-1.0$ $-1.89\pm 0.213$ $-1.68\pm 0.236$ $-1.22\pm 0.228$
-Table 3: Priors and best-fit values for the $\Lambda$CDM and the $w$CDM parameters obtained from synthetic data with $\Delta m=0.2$.
-4.2 Application to Synthetic Data
-For simulations, we first consider the parameters of the Pantheon sample: 1048 supernovae in the redshift range $z\in[0.01,2.26]$ with magnitude errors, $\Delta m\simeq 0.2$ (for simulation, we assume Gaussian random errors, $\mathcal{N}(0,0.20)$).
-We also consider a future SNIa project with errors four times smaller in magnitude, $\mathcal{N}(0,0.05)$.
-This is possible with a sample 16 times larger than the Pantheon sample with the magnitude errors obtained from binning the data in redshift (this might be achieved with Zwicky Transient Facility (ZTF)[30] and the Large Synoptic Sky Survey Telescope (LSST) [31, 32], albeit with higher supernovae density at $z<0.5$, see, e.g.
-[33] for details).
-The cosmological and other parameters are fixed to the following values: $\Omega_{\rm M}=0.30$, $H_{0}=73~{}\rm km~{}s^{-1}~{}Mpc^{-1}$, $w=-1.0$, and the absolute magnitude of supernova, $M=-19.24$.
-To account for peculiar velocities, each supernova is assigned an error on its redshift, $\mathcal{U}(0,0.003)$, drawn from a uniform distribution.
+${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
+$\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.350^{+0.037}_{-0.029}$ $0.330^{+0.051}_{-0.033}$ $0.320^{+0.057}_{-0.034}$
+$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $74.34\pm 0.37$ $74.03\pm 0.58$ $73.83\pm 0.74$
+$w$ $\mathcal{U}(-2,0)$ $-1.24^{+0.15}_{-0.13}$ $-1.16^{+0.18}_{-0.16}$ $-1.11^{+0.20}_{-0.17}$
+Table 2: Priors and best-fit values for the spatially-flat $w$CDM model using Pantheon data.
+The results are shown for three different estimators: $\mathcal{E}_{1}$ ignores peculiar velocities, $\mathcal{E}_{2}$ includes them in the linear approximation, and $\mathcal{E}_{3}$ incorporates them exactly.
+IV.2 Application to Synthetic Data
+Figure 3: Constraints on the $\Lambda$CDM parameters from synthetic data with magnitude error $\sigma_{m}=0.2$ (left) and $\sigma_{m}=0.05$ (right) The colour scheme of Figure 1 is followed.
+As in Figure 2, the errors on the estimated parameters are marginally larger for estimators $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$.
+In addition, the analysis of synthetic data shows that the neglect of peculiar velocities might bias the estimation of cosmological parameters.
+The Figure in conjunction with Figure 1 shows that this bias might not play an important role in the current data (left panel), but it would become important in the interpretation of the future data (right panel).
+[…]
+${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
+$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.325\pm 0.018$ $0.313\pm 0.020$ $0.306^{+0.020}_{-0.022}$
+$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $72.04\pm 0.34$ $72.34\pm 0.41$ $72.55\pm 0.46$
+$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.287^{+0.067}_{-0.049}$ $0.317^{+0.065}_{-0.042}$ $0.331^{+0.061}_{-0.040}$
+$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $71.85\pm 0.46$ $72.48\pm 0.68$ $72.80^{+0.73}_{-0.97}$
+$w$ $\mathcal{U}(-2,0)$ $-1.0$ $-0.92^{+0.16}_{-0.12}$ $-1.05^{+0.21}_{-0.17}$ $-1.13^{+0.25}_{-0.18}$
+Table 3: Priors and best-fit values for the $\Lambda$CDM and $w$CDM parameters obtained from synthetic data with $\sigma_{m}=0.2$.
+Results are shown for three estimators: $\mathcal{E}_{1}$, which ignores peculiar velocities; $\mathcal{E}_{2}$, which uses the linear approximation; and $\mathcal{E}_{3}$, which models them exactly.
+We evaluated the efficacy of our estimators by testing them on synthetic data.
+In order to compare our results for the Pantheon data in the previous section with our results for synthetic data, we adopted the parameters of the Pantheon sample comprising 1048 supernovae and their redshifts that span the redshift range $z\in[0.01,2.26]$.
+For the first case, we fixed the quality of the data at roughly the same level as Pantheon data, assuming the supernova magnitudes to have Gaussian errors with mean zero and standard deviation $\sigma_{m}=0.2$.
+These results can be directly compared with the Pantheon results to determine if they exhibit similar patterns.
+We also consider a future SNIa survey with smaller magnitude errors $\sigma_{m}=0.05$.
+Such small errors are possible with a sample 16 times larger [^5] This data quality might be achieved with the Zwicky Transient Facility (ZTF) Graham et al.
+(2019) and the Large Synoptic Sky Survey Telescope (LSST) Abell et al.
+(2009); Ivezić et al.
+(2019), albeit with higher supernova density at $z<0.5$, see, e.g.Garcia et al.
+(2020) for details).
+than the Pantheon sample with the magnitude errors obtained by binning the data in redshift.
+Note that we used the Pantheon redshifts even for this case to allow for easy comparison.
+Each supernova in this redshift range is assigned a peculiar velocity that is assumed to be a Gaussian random number with zero mean and standard deviation $500\,\rm km\,sec^{-1}$.
+As mentioned before, in order to quantify the effect of reduced errors on magnitude alone, we used the same redshifts as used in the first case.
+Moreover, the random errors used to generate the data are also the same, except for a reduction of a factor of four to achieve $\sigma_{m}=0.05$.
+The cosmological and other parameters are fixed to the following values: $\Omega_{\rm M}=0.30$, $H_{0}=73\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$, $w=-1.0$, and the absolute magnitude of the supernova $M=-19.24$.
 Our main results are shown in Figures 3 and 4 together with Tables 3 and 4.
-A comparison of the analysis of synthetic data with Figures 1 and 2 enables us to discern the following: (a) there is bias in the estimation of cosmological parameters using estimators ${\mathcal{E}}_{1}$ and ${\mathcal{E}}_{2}$.
-The bias is larger if the peculiar velocity contribution is totally neglected (${\mathcal{E}}_{1}$) and is partly alleviated if the linear approximation (Appendix A) is used and (b) while the overlapping confidence regions for $\Delta m=0.2$ do not allow us to determine this bias unambiguously using the current data, the analysis of the future data without taking into account peculiar velocities of host galaxies will result in significant bias in the determination of cosmological parameters.
-In particular, the analysis of SNIa data without accounting for peculiar velocities could bias the nature of dark energy, an important issue from the point of view of both cosmology and theoretical physics, and much debated in recent times following DESI results (e.g.
-[4]).
-Figure 4: Constraints on the $w$CDM parameters from synthetic data with $\Delta m=0.2$ (left) and $\Delta m=0.05$ (right).
-Models Parameters Priors True values Best-fit values
+We carried out many more simulations for our investigations and the results presented here are representative of the observed trends.
+A comparison of our analysis of synthetic data with the input cosmology enables us to discern the following.
+1) There is a random error in the estimation of cosmological parameters using the estimator ${\mathcal{E}}_{1}$.
+The random error changes when synthetic data are drawn with a different set of random errors, and it can either overestimate or underestimate the cosmological parameters.
+The figures show a typical case, but the errors for real data could be larger or smaller.
+2) The estimators ${\mathcal{E}}_{2}$ and ${\mathcal{E}}_{3}$ differ only slightly and are statistically consistent with each other.
+Their agreement with the input values of the cosmological parameters is excellent.
+The linear approximation estimator ${\mathcal{E}}_{2}$ seems adequate for both current and future data.
+3) The discrepancies between ${\mathcal{E}}_{1}$ and the other two estimators, ${\mathcal{E}}_{2}$ and ${\mathcal{E}}_{3}$, become more pronounced with the quality of the data as the magnitude errors of the supernovae decrease from $\sigma_{m}=0.2$ to $\sigma_{m}=0.05$.
+4) The errors in the cosmological parameters increase with the inclusion of redshift errors.
+This is not unexpected since, for variable redshifts, a wider range of cosmological parameters can fit the data.
+5) Compared with the results for $\sigma_{m}=0.2$, the Pantheon results in Figures 1 and 2 show consistent trends.
+This indicates that our assumption for the presence of noise due to peculiar motion of host galaxies in the Pantheon data is reasonable.
+In summary, although the results from the Pantheon data do not significantly improve by taking peculiar velocities into account, we find that analysis of the future data without taking into account peculiar velocities of host galaxies may result in significant bias in the determination of cosmological parameters.
+In particular, the analysis of SNIa data without accounting for peculiar velocities could bias the nature of dark energy, which is an important issue from the point of view of both cosmology and theoretical physics, and has been much debated in recent times following DESI results (e.g.
+Abdul Karim et al.
+(2025)).
+Figure 4: Constraints on the $w$CDM parameters from synthetic data with $\sigma_{m}=0.2$ (left) and $\sigma_{m}=0.05$ (right).
+The colour scheme of Figure 1 is followed.
+As in Figure 1, the errors on the estimated parameters are marginally larger for estimators $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$.
+In addition, the analysis of synthetic data shows that the neglect of peculiar velocities might bias the estimation of cosmological parameters.
+The Figure in conjunction with Figure 2 shows that this bias might not play an important role in the current data (left panel), but it would become important in the interpretation of the future data (right panel).
 […]
-$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.30$ $0.233\pm 0.005$ $0.280\pm 0.005$ $0.298\pm 0.005$
-$H_{0}$ $\mathcal{U}(60,80)$ $73.04$ $75.42\pm 0.19$ $74.03\pm 0.129$ $73.22\pm 0.118$
-$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.2,0.4)$ $0.30$ $0.38\pm 0.010$ $0.322\pm 0.012$ $0.310\pm 0.015$
-$H_{0}$ $\mathcal{U}(60,80)$ $73.04$ $77.74\pm 0.252$ $74.86\pm 0.304$ $73.41\pm 0.243$
-$w$ $\mathcal{U}(-3,0)$ $-1.0$ $-1.87\pm 0.08$ $-1.20\pm 0.07$ $-1.05\pm 0.06$
-Table 4: Priors and best-fit values for the $\Lambda$CDM and the $w$CDM parameters obtained from synthetic data with $\Delta m=0.05$.
-5 Conclusion and Discussion
-Cosmology has become a precision science in the past few decades.
-This has been achieved by an unprecedented increase in both the quality and quantity of data—Planck CMB temperature and polarization maps, Large galaxy surveys such as SDSS, and the detection of thousands of SNIa, etc.
-This also means physical effects hitherto considered beyond the ability of astronomical instruments are well within reach.
-However, such rapid progress also requires one to constantly improve existing statistical techniques and develop new ones.
+${\mathcal{E}_{1}}$ ${\mathcal{E}_{2}}$ ${\mathcal{E}_{3}}$
+$\Lambda$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.314\pm 0.004$ $0.299\pm 0.006$ $0.301^{+0.005}_{-0.006}$
+$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $72.57\pm 0.09$ $72.92\pm 0.15$ $72.88^{+0.16}_{-0.15}$
+$w$CDM $\Omega_{\rm M}$ $\mathcal{U}(0.1,0.5)$ $0.30$ $0.269^{+0.017}_{-0.015}$ $0.320^{+0.018}_{-0.015}$ $0.323^{+0.020}_{-0.016}$
+$H_{0}$ [km/s/Mpc] $\mathcal{U}(60,80)$ $73.0$ $72.31\pm 0.12$ $73.29\pm 0.33$ $73.25^{+0.41}_{-0.46}$
+$w$ $\mathcal{U}(-2,0)$ $-1.0$ $-0.88\pm 0.04$ $-1.09\pm 0.07$ $-1.10^{+0.10}_{-0.08}$
+Table 4: Priors and best-fit values for the $\Lambda$CDM and $w$CDM parameters obtained from synthetic data with $\sigma_{m}=0.05$.
+Results are shown for three estimators: $\mathcal{E}_{1}$, which neglects peculiar velocities; $\mathcal{E}_{2}$, which uses a linear approximation; and $\mathcal{E}_{3}$, which models them exactly.
+V Conclusion and Discussion
+Cosmology has become a precision science in the last few decades.
+This has been achieved by an unprecedented increase in the quality and quantity of data, namely, Planck CMB temperature and polarization maps, large galaxy surveys such as SDSS, the detection of thousands of SNIa, etc.
+This also means that physical effects hitherto considered beyond the ability of astronomical instruments are well within reach.
+However, such rapid progress also requires one to continually improve existing statistical techniques and develop new ones.
 This paper is an effort in that direction.
-The aim of the paper is two-fold.
-First, we develop a general Bayesian method to analyse data with errors in both independent and dependent variables (discussed in detail in section 2.1.1).
-Second, we employ this statistical estimator to SNIa data to test its applicability, efficacy, and relevance to the current and future SNIa and other cosmological data sets.
-In particular, our aim is to discern the role of peculiar velocities of host galaxies of SNIa on the determination of cosmological parameters.
-In Figures 1 and 2, we apply our proposed method on the Pantheon SNIa data for two cosmological models we employ for illustration.
-The figures show that the current data do not require us to account for the nonzero peculiar velocity of the SNIa host galaxies.
-We next apply our estimators on synthetic data, which partly confirms our results with Pantheon data, for parameters of the current data (Figures 3 and 4).
-However, we show that the future data (e.g.
-Zwicky Transient Facility (ZTF)[30] and the Large Synoptic Sky Survey Telescope (LSST) [31, 32]) will be very sensitive to peculiar velocities of SNIa host galaxies, and we would need appropriate statistical tools (section 2.1.1) to extract cosmological parameters from such a data set without bias (Figures 3 and 4).
-One possible improvement of our work is more detailed modelling of peculiar velocities of host galaxies.
+In this paper, we have developed general and approximate Bayesian estimators (discussed in detail in Section II.1.1) to analyze data with errors in both independent and dependent variables and applied them to existing and synthetic SNIa data to test their applicability, efficacy, and relevance to current and future SNIa and other cosmological data sets.
+In particular, we have examined the role of peculiar velocities of host galaxies of SNIa in the determination of cosmological parameters.
+We have applied our method to the Pantheon SNIa data focusing on two cosmological models.
+Our results show that measurement of cosmological parameters with the current data does not change significantly by including peculiar velocities of the SNIa host galaxies.
+However, our results become more consistent with $\Lambda$CDM when peculiar velocities are taken into account, which is significant because of implications for the nature of dark energy.
+Application of our estimators to synthetic data with errors similar to the Pantheon data shows trends consistent with those observed in the results for the Pantheon data, suggesting that our assumptions about peculiar velocities are valid.
+We also showed that the future data with smaller magnitude errors (e.g.
+Zwicky Transient Facility (ZTF)Graham et al.
+(2019) and the Large Synoptic Sky Survey Telescope (LSST) Abell et al.
+(2009); Ivezić et al.
+(2019)) will be much more sensitive to peculiar velocities of host galaxies of SNIa; therefore, the statistical tools developed in this paper will become essential for extracting cosmological parameters from future data sets.
+Our approach differs from previous studies based on N-body simulations and galaxy data (e.g., Gavas et al.
+(2024); Wu and Huterer (2017); Wojtak et al.
+(2014); Peterson et al.
+(2024); Davis et al.
+(2011); Peterson et al.
+(2022); Giani et al.
+(2024)) to assess the impact of peculiar velocities, as it is based only on the existence of the magnitude–redshift relationship, and treats the effect of peculiar velocities as noise in the measurement of supernova redshift.
+Our method does not need to make prior assumptions about the correlation of velocities, e.g., the correlated large-scale component and the random small-scale component as expected in the $\Lambda$CDM model.
+One possible extension of our work is a more detailed modeling of the peculiar velocities of host galaxies.
 It is difficult to predict the peculiar velocities of host galaxies without N-body simulations (e.g.
-[19] for a recent effort on the impact of peculiar velocities on Hubble’s constant measurement).
-This velocity would also depend on the location of host galaxies, e.g., a host galaxy in a rich cluster of galaxies as opposed to an isolated galaxy.
-In addition, and importantly, velocities are correlated across 100s of Mpc (e.g.
-[34] for details and further references), and this correlated component needs to be accounted for in a more detailed discussion (e.g.
-[33]).
-Also, the aim of our paper is to treat peculiar velocities as nuisance parameters and to determine their impact on the interpretation of SNIa data.
+Gavas et al.
+(2024) for a recent effort on the impact of peculiar velocities on Hubble constant measurement).
+This velocity would also depend on the location of host galaxies, e.g.
+a host galaxy in a rich cluster of galaxies as opposed to an isolated galaxy.
+In addition and importantly, velocities are correlated across hundreds of Mpc (see, e.g., Dodelson and Schmidt (2020) for details and further references), and this correlated component needs to be accounted for in a more detailed discussion (e.g., Garcia et al.
+(2020)).
+In addition, the aim of our paper is to treat peculiar velocities as nuisance parameters and to determine their impact on the interpretation of the SNIa data.
 However, it is clear from the context that one might seek to infer cosmological peculiar velocities from the SNIa data.
 From cosmological data, peculiar flows are difficult to measure for individual objects and their presence is revealed through redshift-space distortion of galaxy correlation functions (e.g.
-[34] for details) and kSZ effect of clusters (e.g.
-[17, 18] for further details).
-Our proposed method might complement this ongoing effort to measure cosmological peculiar flows.
+Dodelson and Schmidt (2020) for details) and kSZ effect of clusters (e.g.
+Aghanim et al.
+(2018); Tanimura et al.
+(2021) for further details).
+Our proposed method complements this ongoing effort to measure peculiar cosmological flows.
 In future work, our aim is to extend our work with a better modeling of peculiar velocities along with studying the feasibility of their detection using the SNIa data.
-In addition, a combined analysis using SNIa data sets Pantheon$+$, DESY5, and Union3 might improve our results, which we hope to investigate in a later work.
+In addition, a combined analysis using SNIa data sets Pantheon$+$, DESY5 and Union3 might improve our results, which we hope to investigate in a later work.
 We also expect that our work will find general applicability in the analysis of other cosmological data sets, which we might study in upcoming works.
@@
 UU also thanks Somnath Bharadwaj, Purba Mukherjee, and Avinash Paladi for their helpful comments.
+We acknowledge the use of the HPC cluster of the Raman Research Institute, Bangalore, India.
 Appendix A Linear Approximation
 The problem of fitting errors-in-variable models is quite straightforward for linear models.
-This is because the linear model allows us to marginalize the likelihood over the true regressor analytically, resulting in a very simple likelihood that depends only on the observed values of the variables [25].
+This is because the linear model allows us to marginalize the likelihood over the true regressor analytically, resulting in a very simple likelihood that depends only on the observed values of the variables Press et al.
+(1992).
 Since our model is not linear, we cannot apply it exactly.
@@
 Apparent magnitude as a function of redshift, $z$, is given by
-$m(z)=M+5\;\log\left(\frac{D_{\rm L}(z)}{\text{Mpc}}\right)+25\;,$ (A.1)
+$m(z)=M+5\log\left(\frac{D_{\rm L}(z)}{\text{Mpc}}\right)+25\;,$ (19)
 where $D_{\rm L}(z)$ is the luminosity distance given by,
-$D_{\rm L}(z)=(1+z)\int_{0}^{z}\frac{c\;dz}{H(z)}.$ (A.2)
-Considering errors in the measurement of both the apparent magnitude, $m$, and redshift, $z$, the likelihood can be expressed as
-$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z^{*}_{i})}{\sigma_{m_{i}}}\right)^{2}-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{z_{i}-z^{*}_{i}}{\sigma_{z_{i}}}\right)^{2}$ (A.3)
+$D_{\rm L}(z)=(1+z)\int_{0}^{z}\frac{c\;dz}{H(z)}.$ (20)
+Considering errors in the measurement of both the apparent magnitude, $m$, and redshift, $z$, and assuming the errors to be Gaussian distributed, the likelihood can be expressed as
+$\ln\mathcal{L}\propto-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z^{*}_{i})}{\sigma_{m_{i}}}\right)^{2}-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{z_{i}-z^{*}_{i}}{\sigma_{z_{i}}}\right)^{2}$ (21)
 This likelihood cannot be used to constrain the parameters of the model from the data since we do not have the true redshift values $z_{*}$, only the observed values $z$.
 Taylor expanding the apparent magnitude about the observed redshift and keeping terms up to linear order in redshift, we can write
-$m(z^{*})=m(z)+m^{\prime}(z)(z^{*}-z)$ (A.4)
+$m(z^{*})=m(z)+m^{\prime}(z)(z^{*}-z)$ (22)
 where $m^{\prime}(z)$ is the derivative of $m(z^{*})$ w.r.t. $z^{*}$ evaluated at $z^{*}=z$, and is given by
-$m^{\prime}(z)=\frac{5}{\ln 10}\left[\frac{1}{1+z}+\frac{1}{H(z)\int_{0}^{z}\frac{dz^{*}}{H(z^{*})}}\right]$ (A.5)
-With the above approximation, eq. (A.4), the likelihood takes the form,
-$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z_{i})}{\sigma_{m_{i}}}\right)^{2}-\\ \frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{m^{\prime 2}(z_{i})}{\sigma^{2}_{m_{i}}}+\frac{1}{\sigma^{2}_{z_{i}}}\right)(z^{*}_{i}-z_{i})^{2}-\frac{m^{\prime}(m_{i}-m(z_{i}))}{\sigma^{2}_{m_{i}}}(z^{*}_{i}-z_{i})\right]$
+$m^{\prime}(z)=\frac{5}{\ln 10}\left[\frac{1}{1+z}+\frac{1}{H(z)\int_{0}^{z}\frac{dz^{*}}{H(z^{*})}}\right]$ (23)
+With the above approximation, Eq. (22), the likelihood takes the form,
+$\ln\mathcal{L}\propto-\frac{1}{2}\sum_{i=1}^{N}\left(\frac{m_{i}-m(z_{i})}{\sigma_{m_{i}}}\right)^{2}-\\ \frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{m^{\prime 2}(z_{i})}{\sigma^{2}_{m_{i}}}+\frac{1}{\sigma^{2}_{z_{i}}}\right)(z^{*}_{i}-z_{i})^{2}-\frac{m^{\prime}(m_{i}-m(z_{i}))}{\sigma^{2}_{m_{i}}}(z^{*}_{i}-z_{i})\right]$
 Integrating over the true but unknown redshift $z^{*}$ using the standard Gaussian integral, we obtain the marginalized likelihood,
-$\ln\mathcal{L}=-\frac{1}{2}\sum_{i=1}^{N}\frac{\left(m_{i}-m(z_{i})\right)^{2}}{\sigma^{2}_{m_{i}}+m^{\prime 2}(z_{i})\;\sigma^{2}_{z_{i}}}$ (A.6)
+$\ln\mathcal{L}\propto-\frac{1}{2}\sum_{i=1}^{N}\frac{\left(m_{i}-m(z_{i})\right)^{2}}{\sigma^{2}_{m_{i}}+m^{\prime 2}(z_{i})\;\sigma^{2}_{z_{i}}}$ (24)
 Now, the likelihood depends only on the observed data.
 If we do not know the values of $\sigma_{z_{i}}$, as in our case, we can fix it at some value or make it a parameter.
-In this work, we chose to fix it at $\sigma_{z}=0.003\;$, which corresponds to the peculiar velocity of $1000$ km/s.
+In this work, we chose to fix it at $\sigma_{z}=(\sigma_{v_{p}}/c)(1+z)\;$, with $\sigma_{v_{p}}=1000$ km/s.